### v3

# Towards Efficient Transferable Preemptive Adversarial Defense

Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy because of its sensitivity to inconspicuous perturbations (i.e., adversarial attacks). Attackers may utilize this sensitivity to manipulate predictions. To defend against such attacks, we have devised a proactive strategy for “attacking” the medias before it is attacked by the third party, so that when the protected medias are further attacked, the adversarial perturbations are automatically neutralized. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first to our knowledge effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing proactive defenses against adversarial attacks. The proposed methodology will be made available on GitHub.
Index Terms:
Transferable preemptive adversarial defense, adaptive reversion attack, proactive defense, adversarial attack.
I Introduction
Deep learning technologies have been demonstrated to be sensitive to inconspicuous perturbations (i.e., adversarial attacks) [1]. An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos. For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access (i.e., hack into) the face authentication system [2, 3]. The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [4]. Defending against such attacks is a critical but unsolved challenge. Passive defenses, on the one hand, generally sacrifice classification accuracy for resistance against threats [5, 6, 7, 8, 9]. Proactive defenses, on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications. Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is not transferable to other systems [10, 11]. Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [12, 13].
Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Figure 1. Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party. Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection. In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized. This strategy has three distinct advantages. (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system. (ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated at the classification model) is breached, all inputs are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets four basic requirements:
  • Preserve clean accuracy and improve robust accuracy: Basic requirements for all defenses.
  • Not require ground truths: It is impractical to label each input manually.
  • Provide an efficient defense: Time cost per process must be satisfactory for user interaction.
  • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
As illustrated in Figure 2, the Fast Preemption framework includes three components: a classifier model for labeling inputs, a backbone model for learning features, and the forward-backward (Fore-Back) cascade learning algorithm for computing protective perturbations. The classifier determines the correctness of the optimization direction. The backbone model greatly determines transferability and time cost. However, the best backbone model must be adversarially trained [5, 14], and its clean accuracy is typically sacrificed for robustness against threats. The Fast Preemption framework uses different models for the classifier and backbone, thereby optimizing both tasks. Additionally, as we demonstrate, the state-of-the-art (SOTA) preemptive defense [10] cannot achieve satisfactory time cost for efficient user interaction. We thus developed the Fore-Back cascade learning algorithm combining forward propagation optimization to achieve rapid convergence and backward propagation learning to alleviate overfitting.
We conducted evaluation experiments and ablation studies on two datasets [15, 16] across four models [5, 6, 7, 8], defending against two typical adversarial attacks [17, 18] in three settings. The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [7, 8, 9, 10, 11]. For example, in an end-to-end comparison with the SOTA preemptive defense [10] on the ImageNet dataset [16], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [18] and was 33 times faster.
Shannon insightfully claimed that “one ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [19].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. As illustrated in Figure 3, Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
The results from reversion attack testing demonstrate that the Preemptive Reversion attack can smartly erase protective perturbations when all defense details are compromised (i.e., a white-box situation), whereas blind JPEG distortion and adversarial purification [9] only further distort an image. However, this white-box situation is impractical with the Fast Preemption strategy for real-world applications because the proposed framework enables users to individually customize their defense settings. Apart from a white-box situation, Fast Preemption protection cannot be reversed using existing reversion approaches.
We summarize our contributions as follows:
  • We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module. It is the first to use different models for the classifier and backbone.
  • We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction.
  • We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
Additionally, we will release a new toolbox upon acceptance that enables plug-and-play implementation with multiple options for each component. This toolbox can also integrate benchmark preemptive defenses.
II Related Work
Existing adversarial defenses are divided into three categories: training-time, test-time, and preemptive.
Training-time defenses produce robust deep learning models that classify adversarial examples (AEs) correctly. The earliest mechanisms, e.g., adversarial training [1], incorporated various types of AEs along with their ground-truth labels in the training dataset. To improve the efficiency of training on large-scale datasets, Shafahi et al. [20] devised accelerated adversarial training in which backward-pass computations are reused. This entails curating a large number of specific real-world samples; Sehwag et al. [21] circumvented this problem by using additional data from proxy distributions learned by advanced generative models for adversarial training. Gowal et al. [5] demonstrated that generative models trained solely on the original dataset can be leveraged to artificially increase the training dataset’s size and improve adversarial robustness against $L_{p}$-norm perturbations. Debenedetti et al. [6] proposed using a custom recipe of vision transformers for adversarial training in which weaker data augmentation is used with warm-up in the augmentation intensity. Inspired by denoising diffusion models [22], Wang et al. [23] conducted adversarial training using data generated by an elucidating diffusion model [24]. Dong et al. [25] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration. The current SOTA robust model, devised by Peng et al. [7], follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions.
Training-time defenses have three shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, robust models can be counterattacked. Third, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al. [25], which was designed to circumvent this problem.
Test-time defenses combine detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs. Among recent studies, Cohen et al. [26] devised a k-nearest neighbor model for ranking the most supportive training samples for any given validation example in training a detector. Zuo and Zeng [27] proposed using a binary classifier, erase-and-restore, that uses the intriguing sensitivity of $L_{2}$-norm attacks by randomly erasing certain pixels in an $L_{2}$-norm AE and then restoring them with an inpainting technique. Liang et al. [28] detected AEs by comparing the classification results of a given sample and a denoised version that was generated through adaptive noise reduction combining scalar quantization and a smoothing spatial filter. Most recently, Hickling et al. [29] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification. For instance, Wang et al. [30] distorted images by randomly nullifying pixels. Guo et al. [31] preprocessed inputs with a nondifferentiable transformation. Raff et al. [32] combined a large number of individually weak defenses into a single randomized transformation to build a strong defense. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious. Ren et al. [33] trained a generative model using autoencoder technology [34]. They claimed the existence of an immune space, in which perturbations have fewer adverse effects, and they inactivated adversarial perturbations by restricting them to that subspace. Nie et al. [9] devised the DiffPure benchmark, which uses denoising diffusion models [22]. DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy.
Lastly, Alfarra et al. [35] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer. However, we believe that the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
Preemptive defenses aim to generate REs that replace original images, which are then discarded. These REs are more challenging to manipulate and thus resist adversarial attacks. Salman et al. [36] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples. They were learned using projected gradient descent (PGD) [17] and improved both in-distribution performance and robustness against unforeseen perturbations. Moon et al. [10] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy. However, its time cost is huge, and its transferability is low. Frosio and Kautz [11] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained. It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
Figure 2: Fast Preemption framework. The classifier labels inputs automatically, which improves clean accuracy and thereby improves defense performance. The backbone model learns crucial features, so it must be adversarially trained. The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
III Preemptive Defense Strategy: Fast Preemption
With the preemptive defense strategy, the image is preemptively “attacked” (manipulated for enhancement), protecting it before being attacked. In this section, we present the framework and algorithm of our Fast Preemption strategy.
III-A Problem Definition
Adversarial attacks and preemptive defenses both manipulate the input by adding subtle perturbations. Specifically, let $x$ and $y$ denote the original image and its ground-truth label, respectively, and let $f_{v}$ denote the victim model. Let $x^{\prime}$ denote the manipulated example (AE or RE), formulated as
$x^{\prime}=x+\delta,\ s.t.\ ||\delta||_{p}\leq\epsilon,$ (1)
where $||\cdot||_{p}$ represents the $L_{2}$- or $L_{\infty}$-norm and $\epsilon$ is the setting of the maximum distance. Therefore, adversarial attacks and preemptive defenses aim to solve different optimization problems (as explained below) by changing $\delta$ with the inputs of $x$, $y$, and $f_{v}$.
III-A1 Adversarial Attack
An adversarial attack leads to misclassification due to subtle perturbations. Given an original image $x$ and its ground truth $y$, the adversarial attacker is able to generate an AE by maximizing the cross entropy (CE) loss of $x$ for label $y$, thereby reducing the model’s confidence in correctly classifying the input as its true label. Therefore, the optimization objective of an adversarial attack for misclassification is
[…]
III-A2 Preemptive Defense
A preemptive defense aims to generate an RE that will make Equation (2) less successful. Specifically, with the preemptive defense strategy, “backward propagation optimization” is used to reverse the adversarial attack perturbations [10]. In other words, the optimization of Equation (2) remains, but output $\delta$ is reversed. The backward propagation optimization is formulated as
[…]
In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy. It directly reverses the optimization direction of Equation (2) from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
[…]
Both the forward and backward optimization learning algorithms will be further addressed in Section III-C.
III-B Framework
As illustrated in Figure 2, the Fast Preemption framework includes three components.
Classifier: The correctness of the input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Equations (3) and (4) can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images on the basis of the ground truth can be regarded as using a perfect classifier.
Backbone Model: With Fast Preemption, the input image is iteratively optimized on the basis of feedback from backbone model $f_{b}$, which may differ from victim model $f_{v}$ in Equations (3) and (4). Therefore, due to its complexity, $f_{b}$ determines the transferable clean and robust accuracies and time cost. To achieve better performance, the backbone model $f_{b}$ must be adversarially trained so it can learn more crucial features instead of random pixels. The limitations of adversarially trained models (i.e., training-time defenses), such as the degradation of clean accuracy, are alleviated by the well-performing nonrobust classifier $f_{c}$ in our framework. As a result, the preemptive defense optimizes the inputs towards the correct labels. Therefore, two different models, classifier $f_{c}$ and backbone $f_{b}$, are integrated into the framework.
Fore-Back Cascade Learning Algorithm: To meet our framework, the optimization problems formulated as Equations (3) and (4) can be reformulated as
[…]
respectively, where $f_{c}$ is our classifier and $f_{b}$ is our backbone model. This cascade learning algorithm is described in detail in the next section.
III-C Forward-Backward (Fore-Back) Cascade Learning
Algorithm
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions.
III-C1 Loss Function
From Equations (5) and (6), the loss function of our Fore-Back cascade learning algorithm is the CE loss:
$\ell(x_{i}|x,f_{c},f_{b})=CE(f_{b}(x_{i}),f_{c}(x)),$ (7)
where $x_{i}$ and $x$ denote the output example of the $i$-th iteration and the original input, respectively, and $f_{c}$ and $f_{b}$ are our classifier and backbone models, respectively. The forward optimization reduces $\ell$, while the backward optimization increases it.
III-C2 Preliminary - Backward Propagation
The use of backward propagation for preemptive defense was introduced in the SOTA preemptive defense strategy [10]. At each step of that strategy, a PGD attack [17] is run first as $x_{1}\rightarrow x_{1}^{a}$ (Figure 2), and then the adversarial perturbations are reversed in accordance with the learning rate and thereby forms protective perturbations, as $x_{1}\rightarrow x_{2}$.
Specifically, backward propagation first runs an $I$-iteration $L_{\infty}$-norm PGD attack from $x_{t}$, which is the output of the previous optimization step (forward or backward). In other words, the backward propagation increases $\ell$ via
$x_{t,0}=x_{t},$ (8)
$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size. The final output of Equation (8) is adversarial example $x_{t}^{a}$. Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}-\eta\cdot(x_{t}^{a}-x_{t}))$ (9)
$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))).$
The SOTA preemptive defense strategy [10] conducts only backward propagation learning. However, adversarial attacks can be randomly initialized, so backward propagation requires many steps to converge (minimum ten steps). The Fast Preemption strategy accelerates defense by running forward propagation before running backward propagation learning.
III-C3 Combining Forward and Backward Propagation Learning
We define forward propagation learning as a direct enhancement by moving away from the decision boundary in accordance with the learning rate ($x_{0}\rightarrow x_{1}$ in Figure 2). Therefore, it first reduces $\ell$ via
$x_{t,0}=x_{t},$ (10)
$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
the final output of which is robust example $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}+\eta\cdot(x_{t}^{r}-x_{t}))$ (11)
$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
which is the same as for Equation (9). Therefore, the above derivations can be summarized as
$x_{t,0}=x_{t},$ (12)
$x_{t,i+1}=\left\{\begin{array}[]{rl}\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\rightarrow\\ \prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\leftarrow\end{array}\right.$
$x_{t+1}=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
$s.t.\ ||x_{*}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation (12) indicates that both forward and backward propagation make the image more robust but with different learning directions for CE loss $\ell$.
Whereas backward propagation requires many steps to converge, forward propagation converges quickly. However, forward propagation causes overfitting of the backbone model, so the defense will be less transferable to other victim models. Fast Preemption alleviates this overfitting: $T_{\rightarrow}$ steps of forward propagation are run until there is slight overfitting; then $T_{\leftarrow}$ steps of backward propagation are run to eliminate the overfitting. We evaluated whether the learning is overfitted by evaluating transferability. With this combination of forward and backward propagation, Fast Preemption attains SOTA performance with only three steps (one step $T_{\rightarrow}$, followed by two steps $T_{\leftarrow}$), which dramatically reduces the time cost of RE generation.
III-C4 Averaging $N$ Fast Gradient Sign Method (FGSM) Examples
Forward and backward propagation both require a backbone optimization algorithm, i.e., an adversarial attack algorithm. For example, The Bi-Level optimization [10] used PGD [17] for backward propagation. However, PGD has to run many sequential iterations to compute superior adversarial perturbations, which greatly increases the time cost. With the Fast Preemption strategy, FGSM [1] is used to speed up the computation of adversarial perturbations.
Since FGSM has difficulty obtaining optimal perturbations, with the Fast Preemption strategy, it is run $N$ times with random initialization to produce $N$ FGSM examples, The perturbations are then averaged to obtain the final output. We demonstrate that averaging the perturbations for PGD examples negligibly affects performance (see Section VII-D), whereas with the Fast Preemption strategy, FGSM optimization is effectively improved while efficiency is preserved.
Furthermore, averaging perturbations in the $N$ FGSM examples is faster than processing a single example with iterative PGD. To generate a single RE, iterative PGD has to run each iteration sequentially from a single input, whereas $N$ FGSM examples can be produced by repeating the single input $N$ times and processing the results in parallel in a single iteration. The improvement attained by averaging the perturbations in $N$ FGSM examples is demonstrated in Table VIII.
Algorithm 1 Fast Preemption
1:Original image $x$, classifier $f_{c}$, backbone model $f_{b}$, maximum $L_{\infty}$-norm distance $\epsilon$, learning rate $\eta$, maximum forward steps $T_{\rightarrow}$ and backward steps $T_{\leftarrow}$, number of examples for averaging perturbations $N$.
2:Robust example $x^{r}$.
3:// Forward Propagation
4:$x_{0}=x$
5:for $t<T_{\rightarrow}$ do
6:     for $n=1,...,N$ do
7:         $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
8:         $x_{t,n}=x_{t}+\delta_{n}$
9:         $x_{t,n}^{r}=\prod_{x,\epsilon}(x_{t,n}-\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM decreases $\ell$
10:     end for
11:     $\delta_{t}^{r}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{r}-x_{t}$$\triangleright$ averages perturbations
12:     $x_{t+1}=x_{t}+\eta\cdot\delta_{t}^{r}$$\triangleright$ direct enhancement
13:end for
14:// Backward Propagation
15:$x_{0}=x_{T_{\rightarrow}}$
16:for $t<T_{\leftarrow}$ do
17:     for $n=1,...,N$ do
18:         $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
19:         $x_{t,n}=x_{t}+\delta_{n}$
20:         $x_{t,n}^{a}=\prod_{x,\epsilon}(x_{t,n}+\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM increases $\ell$
21:     end for
22:     $\delta_{t}^{a}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{a}-x_{t}$$\triangleright$ averages perturbations
23:     $x_{t+1}=x_{t}-\eta\cdot\delta_{t}^{a}$$\triangleright$ reverses the attack
24:end for
25:return $x^{r}=x_{T_{\leftarrow}}$
III-C5 Algorithm
The complete Fast Preemption algorithm is presented in Algorithm 1. With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
IV Adaptive Reversion Attack:
Preemptive Reversion
A reversion attack is used to evaluate whether a preemptive defense is reversible. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario. Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
IV-A Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Figure 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Figure 3. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
The complete algorithm is presented in Algorithm 2.
Figure 3: Illustration of Preemptive Reversion algorithm. It first conducts a secondary preemptive defense on RE $x^{r}$ and then reverses the perturbations of the secondary defense: $(x^{r})^{r}-x^{r}$.
Algorithm 2 Preemptive Reversion
1:Robust example $x^{r}$.
2:Reversed example $\tilde{x}$.
3:$(x^{r})^{r}=PreemptiveDefense(x^{r})$$\triangleright$ secondary defense
4:$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r})$$\triangleright$ reverses the secondary defense
5:return $\tilde{x}$
IV-B Attack Scenarios
We define two scenarios for a reversion attack in accordance with the knowledge the attacker has about the backbone model and defense algorithm:
  • White-Box: The attacker knows both the backbone model and defense algorithm (including settings), so the attacker can conduct a secondary defense the same as the user does. This is the strongest, yet impractical, assumption because different users may use different settings.
  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (due to infinite selections for users). The major challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our proposed Fast Preemption defense, but it is feasible only in the white-box scenario. Without full knowledge of the defense, Fast Preemption is reliable.
IV-C Evaluation Protocol
If a reversion attack succeeds, the improved clean accuracy should drop. However, image distortion can create a similar effect. Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack. To address this issue, we proposed a more comprehensive protocol that can distinguish successful reversion from distortion. First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy. In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded. Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
TABLE I: 1,000 original CIFAR-10 [15] images show similar accuracy with that of full test set against AutoAttack (Standard) [18] with training-time defense [7] @ $L_{\infty}\ \epsilon=8/255$.
     Number of Test Samples      Clean (%)      Robust (%)
     10,000 (full)      93.27      71.07
     1,000 (ours)      93.5      71.4
     100      95      68
V Experiments
V-A Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [15] (size of 32$\times$32) and ImageNet [16] (cropped to 224$\times$224 to match the requirements of the classifier and backbone model). CIFAR-10: We conducted ablation studies using 1,000 samples and then tested our defense and compared the results with benchmarks using another 1,000 samples (resembling the validation and test datasets). ResNeXt29-32x4d [37] and PreActResNet-18 [5] were selected as the classifier and backbone model, respectively. Because Fast Preemption does not require model training, 1,000 samples provided a satisfactory diversity and distribution for testing, as demonstrated by the results listed in Table I. ImageNet: We evaluated Fast Preemption and benchmarks using the development dataset from the NIPS 2017 competition, including 1,000 carefully selected ImageNet samples. DeiT Base [38] and ConvNeXt-L+ConvStem [14] were selected as the classifier and backbone model, respectively. The model architectures are listed in Table II.
TABLE II: Model architectures.
       Dataset        Architecture
       CIFAR-10        XCiT-L12 [6]
       RaWideResNet-70-16 [7]
       PreActResNet-18 [5] (backbone)
       ResNeXt29-32x4d [37] (classifier)
       ImageNet        Swin-L [8]
       ConvNeXt-L+ConvStem [14] (backbone)
       DeiT Base [38] (classifier)
V-B Benchmark Adversarial Defenses
Training-time defenses: We compared Fast Preemption with PreActResNet-18 [5], XCiT-L12 [6], and RaWideResNet-70-16 [7] on the CIFAR-10 dataset. We further compared it with Swin-L [8] on the ImageNet dataset. RaWideResNet-70-16 [7] and Swin-L [8] are the SOTA training-time defenses for each dataset, respectively. As listed in Table II, the model architectures of the SOTA training-time defenses differ from the backbone models, so our experiments are a reasonable way to evaluate transferability.
Test-time defenses: We used DiffPure [9] as the benchmark test-time defense. The inputs (regardless of original or malicious) were purified by DiffPure and then fed into RaWideResNet-70-16[7] or Swin-L [8] in accordance with the dataset used.
Preemptive defenses. We compared Fast Preemption with two benchmark preemptive defenses: the SOTA preemptive defense, Bi-Level [10], and the latest preemptive defense, $A^{5}$ [11]. We used the best defense recipe, “RC (robustifier and classifier),” given by Frosio and Kautz [11].
V-C Attacks
Adversarial attacks. We evaluated defenses against two adversarial attacks: PGD [17] and AutoAttack [18]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [39] and Square Attack [40]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [41]).
Reversion attacks. We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [9]).
V-D Evaluation metrics
We used both clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective. Here, accuracy refers to the ratio of predictions that match their ground-truth labels. Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs. Additionally, to capture the image quality, the structural similarity index measure (SSIM) [42] was computed with respect to the original images. A larger SSIM indicates a more subtle change.
V-E Implementation
Our implementation partially leveraged the ART [43] and RobustBench [44] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
VI Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Table III.
TABLE III: Experiment Settings.
Type Approach Settings
Adversarial PGD $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
Attack AutoAttack
Reversion Attack JPEG Quality Factor = 50%
3
DiffPure Diffusion Timestep = $0.1^{1}/0.15^{2}$
3
Ours Same with our defense
Defense DiffPure Diffusion Timestep = $0.1^{1}/0.15^{2}$
3
Bi-Level $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
$\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
3
$A^{5}$ $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
3
Ours $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
  • 1 for CIFAR-10; 2 for ImageNet.
  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
TABLE IV: End-to-end performance against AutoAttack [18] and comparison with benchmark adversarial defenses.
Dataset Defense Time Cost (s) Clean (%) Standard $L_{\infty}$ (%) Standard $L_{2}$ (%) Random $L_{\infty}$ (%)
CIFAR-10 Training-Time RaWideResNet-70-16 [7] 0 93.3 72.8 69.3 73.4
8
Test-Time DiffPure [9] 0.3 91.4 75.8 83.7 76.2
8
Preemptive $A^{5}$ [11] 0.01 93.2 72.4 68.1 73.0
8
Bi-Level [10] 28 92.0 80.1 78.0 80.2
8
Fast Preemption (ours) 0.04 95.3 85.6 84.5 85.7
ImageNet Training-Time Swin-L [8] 0 93.0 67.0 59.5 67.5
8
Test-Time DiffPure [9] 1.2 88.8 73.7 79.4 73.8
8
Preemptive Bi-Level [10] 43 93.6 71.5 62.8 72.3
8
Fast Preemption (ours) 1.3 95.1 77.1 67.4 78
8
Fast Preemption (ours) + DiffPure [9] 2.5 93.3 82.9 87.6 83.4
TABLE V: Preemptive defenses resist reversion attacks.
Original (%) Preemptive Defense Distortion Purification Preemptive Reversion (ours)
9
Classifier Backbone Approach Clean (%) JPEG (%) DiffPure (%) White (%) Black (%)
93.5 Bi-Level Bi-Level Bi-Level 91.9 90.5 90.8 93.1 90.7
9
Bi-Level PreActResNet-18 Bi-Level 91.2 90.6 90.7 93.4 91.0
9
Bi-Level PreActResNet-18 Ours 91.3 90.5 90.5 93.4 90.6
9
ResNeXt29-32x4d PreActResNet-18 Ours 95.5 94.4 94.8 94.0 94.6
  • Numbers in red demonstrate that only our reversion attack is feasible but only in the white-box setting.
Figure 4: Perturbations of adversarial attacks and proposed preemptive defenses. Observation of differences between original image and RE show that Fast Preemption perturbations are unnoticeable. Comparison of perturbations between AutoAttack and Fast Preemption show that preemptive defense is more likely to succeed when perturbations are similar.
VI-A Robustness against Adversarial Attacks
We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [18] and victim models [7, 8]. Table IV compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [18]. Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive. For example, in comparison with the least time costly method RaWideResNet-70-16 [7] with the CIFAR-10 dataset, Fast Preemption had 2% and 12.8% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation. Compared with the SOTA preemptive defense, Bi-Level [10], which had the second best accuracy, Fast Preemption had 3.3% and 5.5% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster. Note that preemptive defenses can work together with training-time and test-time defenses. For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last row of Table IV.
Figure 4 illustrates the results of qualitative analysis. A visual comparison of the original image and RE (average SSIM 0.962) revealed that the perturbations generated by Fast Preemption are unnoticeable. Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones. Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [45]:
$\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
$\delta_{Gray}=$ $\ 0.299\cdot\delta_{RGB}.R+$
$\ 0.587\cdot\delta_{RGB}.G+0.114\cdot\delta_{RGB}.B,$
where $\delta_{RGB}.R$, $\delta_{RGB}.G$, and $\delta_{RGB}.B$ represent the red, green, and blue channels, respectively.
VI-B Resistance against Reversion Attacks
We conducted reversion attacks to evaluate whether preemptive defenses are reversible. Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense. Preemptive Reversion was evaluated in both white- and black-box scenarios (see Section IV-B). The results are summarized in Table V. The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense. The other reversion attacks failed because they further reduced clean accuracy (due to distortion). Nevertheless, Preemptive Reversion is only feasible in the white-box setting. Without full knowledge of the defense, Fast Preemption is irreversible. Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails. However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
Figure 5: Perturbations before and after proposed Preemptive Reversion attack. White-box reversion successfully erased most protective perturbations while black-box reversion failed.
Moreover, the results in the fourth row of Table V, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive. The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion. However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion. Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
Figure 6: Compared with the SOTA preemptive defense Bi-Level [10], our framework improved the clean and robust accuracies, and our learning algorithm reduced the time cost for defense.
VII Ablation Studies
We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset. The attacks were PGD [17] and AutoAttack (Standard) [18] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. As illustrated in Figure 6, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [10] in terms of clean and robust accuracies and time cost. These studies and results were aimed at identifying the optimal strategies.
VII-A Classifier
In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Table VI, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
TABLE VI: Performance with different classifiers.
Ablation Time Cost(s) Transferable Defense
6
Classifier Accuracy(%) Clean(%) PGD(%) AutoAttack(%)
Ground Truths 100 28 97.2 86.9 83.1
ResNeXt 95.9 28 95.4 85.7 82.7
Bi-Level 89.1 29 91.9 83.9 81.0
  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [10]. The transferable defenses were all evaluated using RaWideResNet-70-16 [7] as the victim model.
VII-B Backbone Model
The backbone model determines the defense time cost and performance. As illustrated in Figure 7, different backbone models learn different features. To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Section I, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [13].
Figure 7: Different backbone models learn different features. PreActResNet-18 [5], RaWideResNet-70-16 [7], and Bi-Level [10] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
TABLE VII: Performance of different backbone models.
Ablation Time Cost (s) White-box Defense Transferable Defense
11
Backbone Clean (%) PGD (%) AutoAttack (%) Clean (%) PGD (%) AutoAttack (%) Clean (%) PGD (%) AutoAttack (%)
Bi-Level [10] 89.1 43.9 42.4 28 95.5 84.9 81.8 95.3 72.7 70.8
RaWideResNet-70-16 [7] 93.5 74.5 71.4 178 95.7 91.6 89.9 95.2 78.1 76.2
PreActResNet-18 [5] 88.1 61.8 58.6 26 93.8 85.9 83.3 95.4 79.6 77.3
  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [37] as the classifier model, and 100 steps backward propagation as Bi-Level [10]. The transferable defenses were all evaluated using XCiT-L12 [6] as the victim model, different from the backbone models.
As shown in Table VII, the best adversarially trained model was RaWideResNet-70-16 [7], so it is not surprising that its white-box defense outperformed that of the other backbone models. However, this most robust model failed to guarantee the best transferability, and the time cost was much higher due to its complexity. We thus empirically selected PreActResNet-18 [5] as the backbone model in Fast Preemption as it had the least time cost and best transferability. Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
VII-C Forward-Backward (Fore-Back) Cascade Learning
As observed in Figure 8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0). In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90). Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Figure 9. The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs. F0-B3). Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
Figure 8: Forward propagation (F) converges faster while backward propagation (B) alleviates overfitting. The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively. The numbers in brackets represent robust accuracy against AutoAttack (Standard). “Alternate” represents the running of forward and backward propagation alternatively for 100 steps in total.
Figure 9: By combining forward (F) and backward (B) propagation, Fast Preemption obtains performance comparable to that of the SOTA preemptive defense [10] by using only three steps (F1–B2). The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively. The numbers in brackets represent robust accuracy against AutoAttack (Standard).
VII-D Optimization Algorithm
The results in Table VIII indicate that averaging $N$ FGSM examples results in the best performance with the least time cost. Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
TABLE VIII: Performance for different $N$ and optimization algorithms.
Ablation Time Cost (s) Transferable Defense
6
Optimization $N$ Samples Clean (%) PGD (%) AutoAttack (%)
PGD 1 0.7 95.3 87.5 85.8
FGSM 1 0.04 95.4 86.8 85.4
PGD 20 0.8 95.2 87.7 85.9
FGSM 20 0.04 95.5 87.8 85.9
  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [7] as the victim model.
VIII Conclusion
We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption. We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths. In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses. We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised.
This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field. The toolbox that will be released supports further development and offers a platform for discussing relevant technologies.

### v4

@@
-Towards Efficient Transferable Preemptive Adversarial Defense
+Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense
 Abstract
-Deep learning technology has brought convenience and advanced developments but has become untrustworthy because of its sensitivity to inconspicuous perturbations (i.e., adversarial attacks).
+Deep learning technology has brought convenience and advanced developments but has become untrustworthy due to its sensitivity to adversarial attacks.
 Attackers may utilize this sensitivity to manipulate predictions.
-To defend against such attacks, we have devised a proactive strategy for “attacking” the medias before it is attacked by the third party, so that when the protected medias are further attacked, the adversarial perturbations are automatically neutralized.
+To defend against such attacks, existing anti-adversarial methods typically counteract adversarial perturbations post-attack, while we have devised a proactive strategy that preempts by safeguarding media upfront, effectively neutralizing potential adversarial effects before the third-party attacks occur.
 This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features.
@@
 With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses.
-We have also devised the first to our knowledge effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised.
+We have also devised the first, to our knowledge, effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised.
 This work provides a new direction to developing proactive defenses against adversarial attacks.
-The proposed methodology will be made available on GitHub.
-Index Terms:
-Transferable preemptive adversarial defense, adaptive reversion attack, proactive defense, adversarial attack.
-I Introduction
-Deep learning technologies have been demonstrated to be sensitive to inconspicuous perturbations (i.e., adversarial attacks) [1].
+1 Introduction
+Deep learning technologies have been demonstrated to be sensitive to adversarial attacks [16].
 An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos.
-For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access (i.e., hack into) the face authentication system [2, 3].
+For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access the face authentication system [43].
 The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [13].
 Defending against such attacks is a critical but unsolved challenge.
-Passive defenses, on the one hand, generally sacrifice classification accuracy for resistance against threats [5, 6, 7, 8, 9].
-Proactive defenses, on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications.
-Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is not transferable to other systems [10, 11].
-Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [12, 13].
+Passive defenses [5, 26, 31, 20], on the one hand, trade off classification accuracy for resistance against threats [46], but may be vulnerable to counterattacks [6, 42].
+Proactive defenses [29, 15], on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications.
+Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is untransferable to other systems.
+Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [35].
 Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
-To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Figure 1.
+To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Fig.
+1.
 Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party.
@@
 (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system.
-(ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system.
-(iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated at the classification model) is breached, all inputs are affected.
+(ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system [14, 35].
+(iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated within a classification model) is breached, all inputs are affected.
 In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings.
@@
   • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
-As illustrated in Figure 2, the Fast Preemption framework includes three components: a classifier model for labeling inputs, a backbone model for learning features, and the forward-backward (Fore-Back) cascade learning algorithm for computing protective perturbations.
-The classifier determines the correctness of the optimization direction.
-The backbone model greatly determines transferability and time cost.
-However, the best backbone model must be adversarially trained [5, 14], and its clean accuracy is typically sacrificed for robustness against threats.
-The Fast Preemption framework uses different models for the classifier and backbone, thereby optimizing both tasks.
-Additionally, as we demonstrate, the state-of-the-art (SOTA) preemptive defense [10] cannot achieve satisfactory time cost for efficient user interaction.
-We thus developed the Fore-Back cascade learning algorithm combining forward propagation optimization to achieve rapid convergence and backward propagation learning to alleviate overfitting.
-We conducted evaluation experiments and ablation studies on two datasets [15, 16] across four models [5, 6, 7, 8], defending against two typical adversarial attacks [17, 18] in three settings.
+We propose the Fast Preemption defense includes three components: a classifier model for automatically labeling inputs, a backbone model for learning features (differ from the classifier), and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations.
+We conducted evaluation on two datasets [24, 1] across six models [17, 5, 32, 26, 29, 11], defending against two typical adversarial attacks [28, 9].
 The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [5, 26, 31, 29, 15].
-For example, in an end-to-end comparison with the SOTA preemptive defense [10] on the ImageNet dataset [16], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [18] and was 33 times faster.
+For example, in an end-to-end comparison with the state-of-the-art (SOTA) preemptive defense [29] on the ImageNet dataset [1], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [9] and was 33 times faster.
 Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [39].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible.
-As illustrated in Figure 3, Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense.
+Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense.
 Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier).
@@
 Apart from a white-box situation, Fast Preemption protection cannot be reversed using existing reversion approaches.
-We summarize our contributions as follows:
-  • We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module.
+We summarize our contributions as follows: (i) We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module.
 It is the first to use different models for the classifier and backbone.
-  • We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction.
-  • We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
-Additionally, we will release a new toolbox upon acceptance that enables plug-and-play implementation with multiple options for each component.
-This toolbox can also integrate benchmark preemptive defenses.
-II Related Work
+(ii) We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction.
+(iii) We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
+2 Related Work
+Existing adversarial defenses can be categorized into three primary types: training-time, test-time, and preemptive defenses. Appendix A provides a more detailed review.
+Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels [16, 38, 37, 17, 11, 45, 12, 32, 46, 5, 26, 40].
+However, they have two shortcomings.
+First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness.
+Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy.
+Test-time defenses are divided to detection and purification.
+Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs [7, 47, 25, 20].
+However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks.
+Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
+In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations.
+Earlier studies used image distortion for purification [44, 18, 33].
+Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious [34, 31].
+The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [6, 42].
+Preemptive defenses seek to generate robust examples (REs) that replace the original images, which are subsequently discarded [36, 29, 15].
+These REs are inherently more difficult to manipulate, thereby providing resistance against adversarial attacks.
+However, such approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, lower robustness compared to other defense mechanisms.
+Lastly, Alfarra et al.
+[2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer.
+However, the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
+Figure 2: Fast Preemption framework.
+The classifier labels inputs automatically, which improves clean accuracy and thereby improves defense performance.
+The backbone model learns crucial features, so it must be adversarially trained.
+The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
+3 Preemptive Defense: Fast Preemption
+With the preemptive defense strategy, the image is preemptively “attacked” (manipulated for enhancement), protecting it before being attacked.
+In this section, we present the framework and algorithm of our Fast Preemption strategy.
+3.1 Problem Definition
+Adversarial attacks and preemptive defenses both manipulate inputs by adding subtle perturbations.
+Specifically, let $x$ and $y$ denote the original image and its ground-truth label, respectively, and let $f_{v}$ denote the victim model.
+Let $x^{\prime}$ denote the manipulated example (AE or RE), formulated as
+$x^{\prime}=x+\delta,\ s.t.\ ||\delta||_{p}\leq\epsilon,$ (1)
+where $\delta$ denotes the perturbations, $||\cdot||_{p}$ represents the $L_{2}$- or $L_{\infty}$-norm and $\epsilon$ is the setting of the maximum distance.
+Therefore, adversarial attacks and preemptive defenses aim to solve different optimization problems (as explained below) by changing $\delta$ with the inputs of $x$, $y$, and $f_{v}$.
+3.1.1 Adversarial Attack
+An adversarial attack leads to misclassification due to subtle perturbations.
+Given an original image $x$ and its ground truth $y$, the adversarial attacker is able to generate an AE by maximizing the cross entropy (CE) loss of $x$ for label $y$, thereby reducing the model’s confidence in correctly classifying the input as its true label.
+Therefore, the optimization objective of an adversarial attack for misclassification is
+[…]
+3.1.2 Preemptive Defense
+A preemptive defense aims to generate REs those will make Eq.
+2 less successful.
+With the preemptive defense strategy, “backward propagation optimization” is used to reverse the adversarial attack perturbations [29].
+In other words, the optimization of Eq.
+2 remains, but output $\delta$ is reversed.
+The backward propagation optimization is formulated as
+[…]
+In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy.
+It directly reverses the optimization direction of Eq.
+2 from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
+[…]
+3.2 Framework
+As illustrated in Fig. 2, the Fast Preemption framework includes three components.
+Classifier: Correct input labels determines the correct optimization direction for defense.
+However, for efficient user interaction, it is impractical to label each input manually.
+Classifier $f_{c}$ labels these inputs automatically.
+Therefore, $y$ in Eq.
+3 and Eq.
+4 can be replaced by $f_{c}(x)$.
+The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy.
+Processing images with ground truths can be regarded as using a perfect classifier.
+Backbone Model: With Fast Preemption, the input image is iteratively optimized on backbone model $f_{b}$, which may differ from victim model $f_{v}$ in Eq.
+3 and Eq.
+4.
+Therefore, due to its complexity, $f_{b}$ determines the transferable clean and robust accuracies and time cost.
+To achieve better performance, the backbone model $f_{b}$ must be adversarially trained so it can learn more crucial features instead of random pixels.
+The limitations of adversarially trained models (i.e., training-time defenses), such as the degradation of clean accuracy, are alleviated by the well-performing nonrobust classifier $f_{c}$ in our framework.
+As a result, the preemptive defense optimizes the inputs towards the correct labels.
+Therefore, two different models, classifier $f_{c}$ and backbone $f_{b}$, are integrated into the framework.
+Fore-Back Cascade Learning Algorithm: To meet our framework, the optimization problems formulated as Eq. 3 and Eq. 4 can be reformulated as
+[…]
+[…]
+respectively, where $f_{c}$ is our classifier and $f_{b}$ is our backbone model. This cascade learning algorithm is described in detail in the next section.
+3.3 Forward-Backward Cascade Learning
+The Fore-Back cascade learning algorithm is used to compute protective perturbations.
+The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth.
+We define this as forward propagation ($\rightarrow$).
+Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models.
+Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation.
+Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability.
+In summary, the forward and backward propagation both enhance robustness, but in different optimization directions.
+The complete algorithm is presented in Algorithm 1 of Appendix B.
+3.3.1 Loss Function
+From Eq. 5 and Eq. 6, the loss function of our Fore-Back cascade learning algorithm is the CE loss:
+$\ell(x_{i}|x,f_{c},f_{b})=CE(f_{b}(x_{i}),f_{c}(x)),$ (7)
+where $x_{i}$ and $x$ denote the output example of the $i$-th iteration and the original input, respectively, and $f_{c}$ and $f_{b}$ are our classifier and backbone models, respectively.
+The forward optimization reduces $\ell$, while the backward optimization increases it.
+3.3.2 Preliminary - Backward Propagation
+The use of backward propagation for preemptive defense was introduced in the SOTA preemptive defense strategy [29].
+At each step of that strategy, a projected gradient descent (PGD) attack [28] is run first as $x_{1}\rightarrow x_{1}^{a}$ (Fig.
+2), and then the adversarial perturbations are reversed in accordance with the learning rate and thereby forms protective perturbations, as $x_{1}\rightarrow x_{2}$.
+Specifically, backward propagation first runs an $I$-iteration $L_{\infty}$-norm PGD attack from $x_{t}$, which is the output of the previous optimization step (forward or backward).
+In other words, the backward propagation increases $\ell$ via
+$x_{t,0}=x_{t},$ (8)
+$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
+$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
+where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size.
+The final output of Eq.
+8 is AE $x_{t}^{a}$.
+Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
+$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}-\eta\cdot(x_{t}^{a}-x_{t}))$ (9)
+$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))).$
+The SOTA preemptive defense strategy [29] conducts only backward propagation learning.
+However, adversarial attacks can be randomly initialized, so backward propagation requires many steps to converge (minimum ten steps).
+The Fast Preemption strategy accelerates defense by running forward propagation before running backward propagation learning.
+3.3.3 Combining Forward and Backward Propagation
+We define forward propagation learning as a direct enhancement by moving away from the decision boundary in accordance with the learning rate ($x_{0}\rightarrow x_{1}$ in Fig.
+2).
+Therefore, it first reduces $\ell$ via
+$x_{t,0}=x_{t},$ (10)
+$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
+$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
+the final output of which is robust example $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
+$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}+\eta\cdot(x_{t}^{r}-x_{t}))$ (11)
+$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
+which is the same as for Eq. 9. Therefore, the above derivations can be summarized as
+$x_{t,0}=x_{t},$ (12)
+$x_{t,i+1}=\left\{\begin{array}[]{rl}\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\rightarrow\\ \prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\leftarrow\end{array}\right.$
+$x_{t+1}=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
+$s.t.\ ||x_{*}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
+where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
+Whereas backward propagation requires many steps to converge, forward propagation converges quickly.
+However, forward propagation causes overfitting of the backbone model, so the defense will be less transferable to other victim models.
+Fast Preemption alleviates this overfitting: $T_{\rightarrow}$ steps of forward propagation are run until there is slight overfitting; then $T_{\leftarrow}$ steps of backward propagation are run to eliminate the overfitting.
+We evaluated whether the learning is overfitted by evaluating transferability.
+With this combination of forward and backward propagation, Fast Preemption attains SOTA performance with only three steps (one step $T_{\rightarrow}$, followed by two steps $T_{\leftarrow}$), which dramatically reduces the time cost of RE generation.
+3.3.4 Averaging $N$ Examples
+Forward and backward propagation both require a backbone optimization algorithm, i.e., an adversarial attack algorithm.
+For example, The Bi-Level optimization [29] used PGD [28] for backward propagation.
+However, PGD has to run many sequential iterations to compute superior adversarial perturbations, which greatly increases the time cost.
+With the Fast Preemption strategy, Fast Gradient Sign Method (FGSM) [16] is used to speed up the computation of adversarial perturbations.
+Since FGSM has difficulty obtaining optimal perturbations, with the Fast Preemption strategy, it is run $N$ times with random initialization to produce $N$ FGSM examples, The perturbations are then averaged to obtain the final output.
+We demonstrate that averaging the perturbations for PGD examples negligibly affects performance, whereas with the Fast Preemption strategy, FGSM optimization is effectively improved while efficiency is preserved.
+Furthermore, averaging perturbations in the $N$ FGSM examples is faster than processing a single example with iterative PGD.
+To generate a single RE, iterative PGD has to run each iteration sequentially from a single input, whereas $N$ FGSM examples can be produced by repeating the single input $N$ times and processing the results in parallel.
+4 Adaptive Attack: Preemptive Reversion
+A reversion attack is used to evaluate whether a preemptive defense is reversible.
+In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks.
+Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario.
+Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
+4.1 Algorithm
+When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig.
+5).
+On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig.
+3.
+The complete algorithm is presented in Algorithm 2 of Appendix B.
+Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$.
+Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
+$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
+Figure 3: Illustration of Preemptive Reversion algorithm.
+It first conducts a secondary preemptive defense on RE $x^{r}$ and then reverses the perturbations of the secondary defense: $(x^{r})^{r}-x^{r}$.
+4.2 Attack Scenarios
+We define two scenarios for a reversion attack in accordance with the knowledge the attacker has about the backbone model and defense algorithm:
+  • White-Box: The attacker knows both the backbone model and defense algorithm (including settings), so the attacker can conduct a secondary defense the same as the user does.
+This is the strongest, yet impractical, assumption because different users may use different settings.
+  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (infinite selections for users).
+The challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
+Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our Fast Preemption defense, but it is feasible only in the white-box scenario.
+Without full knowledge of the defense, Fast Preemption is reliable.
+4.3 Evaluation Protocol
+If a reversion attack succeeds, the improved clean accuracy should drop.
+However, image distortion can create a similar effect.
+Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack.
+To address this issue, we proposed a comprehensive protocol that can distinguish successful reversion from distortion.
+First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy.
+In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded.
+Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
+5 Experiments
+5.1 Datasets, Classifier, and Backbone Models
+We conducted experiments on two datasets: CIFAR-10 [24] and ImageNet [1].
+For CIFAR-10, ResNeXt29-32x4d [19] and PreActResNet-18 [17] were selected as the classifier and backbone model, respectively.
+For ImageNet, DeiT Base [41] and ConvNeXt-L+ConvStem [40] were selected as the classifier and backbone model, respectively.
+The model architectures are listed in Tab.
+3 of Appendix C.
+5.2 Attacks and Defenses
+Adversarial attacks.
+We evaluated defenses against two adversarial attacks: PGD [28] and AutoAttack [9].
+For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [8] and Square Attack [3]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [4]).
+Reversion attacks.
+We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective.
+Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [31]).
+Training-time defenses: We compared Fast Preemption with WideResNet-94-16 [5] and Swin-L [26] on the CIFAR-10 and ImageNet datasets, respectively.
+The model architectures of these defenses differ from the backbone models.
+Test-time defenses: We used DiffPure [31] as the benchmark test-time defense.
+The inputs were purified by DiffPure and then fed into WideResNet-94-16 [5] or Swin-L [26] in accordance with the dataset used.
+Preemptive defenses.
+We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [29], and the latest preemptive defense, $A^{5}$ [15].
+5.3 Evaluation metrics
+We used clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective.
+Here, accuracy refers to the ratio of predictions that match their ground-truth labels.
+Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs.
+Table 1: End-to-end performance against AutoAttack [9] and comparison with benchmark adversarial defenses.
+Dataset Defense Time Cost (s) Clean (%) Standard $L_{\infty}$ (%) Standard $L_{2}$ (%) Random $L_{\infty}$ (%)
+CIFAR-10 Training-Time WideResNet-94-16 [5] 0 93.4 74.6 69.6 74.8
+8
+Test-Time DiffPure [31] 0.3 91.4 78.1 84.9 78.4
+8
+Preemptive $A^{5}$ [15] 0.01 93.5 73.8 69.0 74.3
+Bi-Level [29] 28 92.5 81.4 78.6 81.6
+Fast Preemption (ours) 0.04 95.0 86.5 84.3 86.7
+8
+Fast Preemption (ours) + DiffPure [31] 0.34 94.2 89.1 92.8 89.7
+ImageNet Training-Time Swin-L [26] 0 93.0 67.0 59.5 67.5
+8
+Test-Time DiffPure [31] 1.2 88.8 73.7 79.4 73.8
+8
+Preemptive Bi-Level [29] 43 93.6 71.5 62.8 72.3
+Fast Preemption (ours) 1.3 95.1 77.1 67.4 78
+8
+Fast Preemption (ours) + DiffPure [31] 2.5 93.3 82.9 87.6 83.4
+  • Bold indicates the best performance.
+6 Performance
+In this section, we present the end-to-end evaluation results and comparisons with benchmarks.
+The results were obtained in the settings listed in Tab.
+4 of Appendix D.
+The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
+Table 2: Preemptive defenses resist reversion attacks.
+Original (%) Preemptive Defense Distortion Purification Preemptive Reversion (ours)
+Classifier Backbone Approach Clean (%) JPEG (%) DiffPure (%) White (%) Black (%)
+93.5 Bi-Level Bi-Level Bi-Level 91.9 90.5 90.8 93.1 90.7
+Bi-Level PreActResNet-18 Bi-Level 91.2 90.6 90.7 93.4 91.0
+Bi-Level PreActResNet-18 Ours 91.3 90.5 90.5 93.4 90.6
+ResNeXt29-32x4d PreActResNet-18 Ours 95.5 94.4 94.8 94.0 94.6
+  • Numbers in red demonstrate that only our reversion attack is feasible but only in the white-box setting.
+Figure 4: Perturbations of adversarial attacks and proposed preemptive defenses.
+Observation of differences between original image and RE show that Fast Preemption perturbations are unnoticeable.
+Comparison of perturbations between AutoAttack and Fast Preemption show that preemptive defense is more likely to succeed when perturbations are similar.
+6.1 Robustness against Adversarial Attacks
+We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [9] and victim models [5, 26].
+Table 1 compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [9].
+Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive.
+For example, in comparison with the least time costly method WideResNet-94-16 [5] with the CIFAR-10 dataset, Fast Preemption had 1.6% and 11.9% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation.
+Compared with the SOTA preemptive defense, Bi-Level [29], which had the second best accuracy, Fast Preemption had 2.5% and 5.1% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster.
+Note that preemptive defenses can work together with training-time and test-time defenses.
+For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last rows of CIFAR-10 and ImageNet in Tab.
+1, respectively.
+Figure 4 illustrates the results of qualitative analysis.
+A visual comparison of the original image and RE revealed that the perturbations generated by Fast Preemption are unnoticeable.
+Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones.
+Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [27]:
+$\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
+$\delta_{Gray}=$ $\ 0.299\cdot\delta_{RGB}.R+$
+$\ 0.587\cdot\delta_{RGB}.G+0.114\cdot\delta_{RGB}.B,$
+where $\delta_{RGB}.R$, $\delta_{RGB}.G$, and $\delta_{RGB}.B$ represent the red, green, and blue channels, respectively.
+6.2 Resistance against Reversion Attacks
+We conducted reversion attacks to evaluate whether preemptive defenses are reversible.
+Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense.
+Preemptive Reversion was evaluated in both white- and black-box scenarios (see Sec.
+4.2).
+The results are summarized in Tab.
+2.
+The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense.
+The other reversion attacks failed because they further reduced clean accuracy (due to distortion).
+Nevertheless, Preemptive Reversion is only feasible in the white-box setting.
+Without full knowledge of the defense, Fast Preemption is irreversible.
+Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails.
+However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
+Figure 5: Perturbations before and after proposed Preemptive Reversion attack. White-box reversion successfully erased most protective perturbations while black-box reversion failed.
+Moreover, the results in the fourth row of Tab.
+2, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive.
+The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion.
+However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion.
+Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
+7 Ablation Studies
+We conducted ablation studies on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset.
+We used AutoAttack (Standard) [9] against a black-box target model RaWideResNet-70-16 [32], employing $L_{\infty}$-norm perturbations with $\epsilon=8/255$.
+As demonstrated in Fig.
+6, Fast Preemption significantly outperforms the SOTA Bi-Level method [29], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to just 0.04 seconds.
+These studies were conducted to identify the optimal configurations, with more extensive ablation results presented in Appendix F.
+Figure 6: Compared with the SOTA preemptive defense Bi-Level [29], our framework improved the clean and robust accuracies, and our learning algorithm reduced the time cost for defense.
+8 Conclusion
+We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption.
+We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths.
+In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses.
+We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised.
+This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field.
+Acknowledgments
+This work was supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grants JPMJCR18A6 and JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3, by JST K Program Grant JPMJKP24C2 Japan, and by the project for the development and demonstration of countermeasures against disinformation and misinformation on the Internet with the Ministry of Internal Affairs and Communications of Japan.
+Supplementary Material
+This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix C), and the experimental configurations for both attacks and defenses (Appendix D).
+Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
+Appendix A Related Work
 Existing adversarial defenses are divided into three categories: training-time, test-time, and preemptive.
-Training-time defenses produce robust deep learning models that classify adversarial examples (AEs) correctly.
+Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels.
 The earliest mechanisms, e.g., adversarial training [16], incorporated various types of AEs along with their ground-truth labels in the training dataset.
+Zhang et al.
+[46] proposed TRADES, which introduces a theoretically grounded approach to balance robustness and accuracy in adversarial training by incorporating a custom loss function that controls the trade-off between clean accuracy and adversarial robustness.
 To improve the efficiency of training on large-scale datasets, Shafahi et al.
@@
 [12] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration.
-The current SOTA robust model, devised by Peng et al.
-[7], follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions.
-Training-time defenses have three shortcomings.
+Peng et al.
+[32] devised a robust model which follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions.
+Bartoldson et al.
+[5] developed the current state-of-the-art (SOTA) robust model by scaling laws for adversarial training to assess robustness in image classifiers, revealing that even with significant computational resources, robustness asymptotes below human-level performance due to flaws in attack formulations generating invalid data, suggesting future efforts should address both computational efficiency and image validity.
+Training-time defenses have two shortcomings.
 First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness.
-Second, robust models can be counterattacked.
-Third, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al.
+Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al.
 [12], which was designed to circumvent this problem.
-Test-time defenses combine detection and purification.
+Test-time defenses are divided to detection and purification.
 Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs.
@@
 [20] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents.
-However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks.
+However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks [6].
 Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
@@
 DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process.
-The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy.
+The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [42].
 Lastly, Alfarra et al.
@@
 However, we believe that the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
-Preemptive defenses aim to generate REs that replace original images, which are then discarded.
-These REs are more challenging to manipulate and thus resist adversarial attacks.
-Salman et al.
-[36] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples.
-They were learned using projected gradient descent (PGD) [17] and improved both in-distribution performance and robustness against unforeseen perturbations.
-Moon et al.
-[10] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy.
-However, its time cost is huge, and its transferability is low.
-Frosio and Kautz [11] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained.
-It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
-Figure 2: Fast Preemption framework.
-The classifier labels inputs automatically, which improves clean accuracy and thereby improves defense performance.
-The backbone model learns crucial features, so it must be adversarially trained.
-The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
-III Preemptive Defense Strategy: Fast Preemption
-With the preemptive defense strategy, the image is preemptively “attacked” (manipulated for enhancement), protecting it before being attacked.
-In this section, we present the framework and algorithm of our Fast Preemption strategy.
-III-A Problem Definition
-Adversarial attacks and preemptive defenses both manipulate the input by adding subtle perturbations.
-Specifically, let $x$ and $y$ denote the original image and its ground-truth label, respectively, and let $f_{v}$ denote the victim model.
-Let $x^{\prime}$ denote the manipulated example (AE or RE), formulated as
-$x^{\prime}=x+\delta,\ s.t.\ ||\delta||_{p}\leq\epsilon,$ (1)
-where $||\cdot||_{p}$ represents the $L_{2}$- or $L_{\infty}$-norm and $\epsilon$ is the setting of the maximum distance.
-Therefore, adversarial attacks and preemptive defenses aim to solve different optimization problems (as explained below) by changing $\delta$ with the inputs of $x$, $y$, and $f_{v}$.
-III-A1 Adversarial Attack
-An adversarial attack leads to misclassification due to subtle perturbations.
-Given an original image $x$ and its ground truth $y$, the adversarial attacker is able to generate an AE by maximizing the cross entropy (CE) loss of $x$ for label $y$, thereby reducing the model’s confidence in correctly classifying the input as its true label.
-Therefore, the optimization objective of an adversarial attack for misclassification is
-[…]
-III-A2 Preemptive Defense
-A preemptive defense aims to generate an RE that will make Equation (2) less successful.
-Specifically, with the preemptive defense strategy, “backward propagation optimization” is used to reverse the adversarial attack perturbations [10].
-In other words, the optimization of Equation (2) remains, but output $\delta$ is reversed.
-The backward propagation optimization is formulated as
-[…]
-In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy.
-It directly reverses the optimization direction of Equation (2) from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
-[…]
-Both the forward and backward optimization learning algorithms will be further addressed in Section III-C.
-III-B Framework
-As illustrated in Figure 2, the Fast Preemption framework includes three components.
-Classifier: The correctness of the input labels determines the correct optimization direction for defense.
-However, for efficient user interaction, it is impractical to label each input manually.
-Classifier $f_{c}$ labels these inputs automatically.
-Therefore, $y$ in Equations (3) and (4) can be replaced by $f_{c}(x)$.
-The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy.
-Processing images on the basis of the ground truth can be regarded as using a perfect classifier.
-Backbone Model: With Fast Preemption, the input image is iteratively optimized on the basis of feedback from backbone model $f_{b}$, which may differ from victim model $f_{v}$ in Equations (3) and (4).
-Therefore, due to its complexity, $f_{b}$ determines the transferable clean and robust accuracies and time cost.
-To achieve better performance, the backbone model $f_{b}$ must be adversarially trained so it can learn more crucial features instead of random pixels.
-The limitations of adversarially trained models (i.e., training-time defenses), such as the degradation of clean accuracy, are alleviated by the well-performing nonrobust classifier $f_{c}$ in our framework.
-As a result, the preemptive defense optimizes the inputs towards the correct labels.
-Therefore, two different models, classifier $f_{c}$ and backbone $f_{b}$, are integrated into the framework.
-Fore-Back Cascade Learning Algorithm: To meet our framework, the optimization problems formulated as Equations (3) and (4) can be reformulated as
-[…]
-[…]
-respectively, where $f_{c}$ is our classifier and $f_{b}$ is our backbone model. This cascade learning algorithm is described in detail in the next section.
-III-C Forward-Backward (Fore-Back) Cascade Learning
-Algorithm
-The Fore-Back cascade learning algorithm is used to compute protective perturbations.
-The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth.
-We define this as forward propagation ($\rightarrow$).
-Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models.
-Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation.
-Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability.
-In summary, the forward and backward propagation both enhance robustness, but in different optimization directions.
-III-C1 Loss Function
-From Equations (5) and (6), the loss function of our Fore-Back cascade learning algorithm is the CE loss:
-$\ell(x_{i}|x,f_{c},f_{b})=CE(f_{b}(x_{i}),f_{c}(x)),$ (7)
-where $x_{i}$ and $x$ denote the output example of the $i$-th iteration and the original input, respectively, and $f_{c}$ and $f_{b}$ are our classifier and backbone models, respectively.
-The forward optimization reduces $\ell$, while the backward optimization increases it.
-III-C2 Preliminary - Backward Propagation
-The use of backward propagation for preemptive defense was introduced in the SOTA preemptive defense strategy [10].
-At each step of that strategy, a PGD attack [17] is run first as $x_{1}\rightarrow x_{1}^{a}$ (Figure 2), and then the adversarial perturbations are reversed in accordance with the learning rate and thereby forms protective perturbations, as $x_{1}\rightarrow x_{2}$.
-Specifically, backward propagation first runs an $I$-iteration $L_{\infty}$-norm PGD attack from $x_{t}$, which is the output of the previous optimization step (forward or backward).
-In other words, the backward propagation increases $\ell$ via
-$x_{t,0}=x_{t},$ (8)
-$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
-$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
-where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size.
-The final output of Equation (8) is adversarial example $x_{t}^{a}$.
-Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
-$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}-\eta\cdot(x_{t}^{a}-x_{t}))$ (9)
-$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))).$
-The SOTA preemptive defense strategy [10] conducts only backward propagation learning.
-However, adversarial attacks can be randomly initialized, so backward propagation requires many steps to converge (minimum ten steps).
-The Fast Preemption strategy accelerates defense by running forward propagation before running backward propagation learning.
-III-C3 Combining Forward and Backward Propagation Learning
-We define forward propagation learning as a direct enhancement by moving away from the decision boundary in accordance with the learning rate ($x_{0}\rightarrow x_{1}$ in Figure 2).
-Therefore, it first reduces $\ell$ via
-$x_{t,0}=x_{t},$ (10)
-$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
-$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
-the final output of which is robust example $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
-$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}+\eta\cdot(x_{t}^{r}-x_{t}))$ (11)
-$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
-which is the same as for Equation (9). Therefore, the above derivations can be summarized as
-$x_{t,0}=x_{t},$ (12)
-$x_{t,i+1}=\left\{\begin{array}[]{rl}\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\rightarrow\\ \prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\leftarrow\end{array}\right.$
-$x_{t+1}=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
-$s.t.\ ||x_{*}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
-where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation (12) indicates that both forward and backward propagation make the image more robust but with different learning directions for CE loss $\ell$.
-Whereas backward propagation requires many steps to converge, forward propagation converges quickly.
-However, forward propagation causes overfitting of the backbone model, so the defense will be less transferable to other victim models.
-Fast Preemption alleviates this overfitting: $T_{\rightarrow}$ steps of forward propagation are run until there is slight overfitting; then $T_{\leftarrow}$ steps of backward propagation are run to eliminate the overfitting.
-We evaluated whether the learning is overfitted by evaluating transferability.
-With this combination of forward and backward propagation, Fast Preemption attains SOTA performance with only three steps (one step $T_{\rightarrow}$, followed by two steps $T_{\leftarrow}$), which dramatically reduces the time cost of RE generation.
-III-C4 Averaging $N$ Fast Gradient Sign Method (FGSM) Examples
-Forward and backward propagation both require a backbone optimization algorithm, i.e., an adversarial attack algorithm.
-For example, The Bi-Level optimization [10] used PGD [17] for backward propagation.
-However, PGD has to run many sequential iterations to compute superior adversarial perturbations, which greatly increases the time cost.
-With the Fast Preemption strategy, FGSM [1] is used to speed up the computation of adversarial perturbations.
-Since FGSM has difficulty obtaining optimal perturbations, with the Fast Preemption strategy, it is run $N$ times with random initialization to produce $N$ FGSM examples, The perturbations are then averaged to obtain the final output.
-We demonstrate that averaging the perturbations for PGD examples negligibly affects performance (see Section VII-D), whereas with the Fast Preemption strategy, FGSM optimization is effectively improved while efficiency is preserved.
-Furthermore, averaging perturbations in the $N$ FGSM examples is faster than processing a single example with iterative PGD.
-To generate a single RE, iterative PGD has to run each iteration sequentially from a single input, whereas $N$ FGSM examples can be produced by repeating the single input $N$ times and processing the results in parallel in a single iteration.
-The improvement attained by averaging the perturbations in $N$ FGSM examples is demonstrated in Table VIII.
 Algorithm 1 Fast Preemption
@@
 25:return $x^{r}=x_{T_{\leftarrow}}$
-III-C5 Algorithm
-The complete Fast Preemption algorithm is presented in Algorithm 1. With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
-IV Adaptive Reversion Attack:
-Preemptive Reversion
-A reversion attack is used to evaluate whether a preemptive defense is reversible.
-In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks.
-Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario.
-Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
-IV-A Algorithm
-When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Figure 5).
-On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Figure 3.
-Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$.
-Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
-$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
-The complete algorithm is presented in Algorithm 2.
-Figure 3: Illustration of Preemptive Reversion algorithm.
-It first conducts a secondary preemptive defense on RE $x^{r}$ and then reverses the perturbations of the secondary defense: $(x^{r})^{r}-x^{r}$.
 Algorithm 2 Preemptive Reversion
@@
 5:return $\tilde{x}$
-IV-B Attack Scenarios
-We define two scenarios for a reversion attack in accordance with the knowledge the attacker has about the backbone model and defense algorithm:
-  • White-Box: The attacker knows both the backbone model and defense algorithm (including settings), so the attacker can conduct a secondary defense the same as the user does.
-This is the strongest, yet impractical, assumption because different users may use different settings.
-  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (due to infinite selections for users).
-The major challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
-Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our proposed Fast Preemption defense, but it is feasible only in the white-box scenario.
-Without full knowledge of the defense, Fast Preemption is reliable.
-IV-C Evaluation Protocol
-If a reversion attack succeeds, the improved clean accuracy should drop.
-However, image distortion can create a similar effect.
-Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack.
-To address this issue, we proposed a more comprehensive protocol that can distinguish successful reversion from distortion.
-First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy.
-In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded.
-Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
-TABLE I: 1,000 original CIFAR-10 [15] images show similar accuracy with that of full test set against AutoAttack (Standard) [18] with training-time defense [7] @ $L_{\infty}\ \epsilon=8/255$.
-     Number of Test Samples      Clean (%)      Robust (%)
-     10,000 (full)      93.27      71.07
-     1,000 (ours)      93.5      71.4
-     100      95      68
-V Experiments
-V-A Datasets, Classifier, and Backbone Models
-We conducted experiments on two datasets: CIFAR-10 [15] (size of 32$\times$32) and ImageNet [16] (cropped to 224$\times$224 to match the requirements of the classifier and backbone model).
-CIFAR-10: We conducted ablation studies using 1,000 samples and then tested our defense and compared the results with benchmarks using another 1,000 samples (resembling the validation and test datasets).
-ResNeXt29-32x4d [37] and PreActResNet-18 [5] were selected as the classifier and backbone model, respectively.
-Because Fast Preemption does not require model training, 1,000 samples provided a satisfactory diversity and distribution for testing, as demonstrated by the results listed in Table I.
-ImageNet: We evaluated Fast Preemption and benchmarks using the development dataset from the NIPS 2017 competition, including 1,000 carefully selected ImageNet samples.
-DeiT Base [38] and ConvNeXt-L+ConvStem [14] were selected as the classifier and backbone model, respectively.
-The model architectures are listed in Table II.
-TABLE II: Model architectures.
+Preemptive defenses aim to generate REs that replace original images, which are then discarded.
+These REs are more challenging to manipulate and thus resist adversarial attacks.
+Salman et al.
+[36] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples.
+They were learned using projected gradient descent (PGD) [28] and improved both in-distribution performance and robustness against unforeseen perturbations.
+Moon et al.
+[29] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy.
+However, its time cost is huge, and its transferability is low.
+Frosio and Kautz [15] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained.
+It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
+Appendix B Algorithms of Fast Preemption and
+Preemptive Reversion
+The complete defense algorithm of Fast Preemption is presented in Algorithm 1.
+With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
+The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
+Appendix C Model Architectures
+We evaluated attacks and defenses using models with various architectures, as listed in Tab. 3.
+Table 3: Model architectures.
       Dataset       Architecture
-       CIFAR-10        XCiT-L12 [6]
+      CIFAR-10       WideResNet-94-16 [5]
       RaWideResNet-70-16 [32]
+      WideResNet-34-20 [29]
+      XCiT-L12 [11]
       PreActResNet-18 [17] (backbone)
@@
       DeiT Base [41] (classifier)
-V-B Benchmark Adversarial Defenses
-Training-time defenses: We compared Fast Preemption with PreActResNet-18 [5], XCiT-L12 [6], and RaWideResNet-70-16 [7] on the CIFAR-10 dataset.
-We further compared it with Swin-L [8] on the ImageNet dataset.
-RaWideResNet-70-16 [7] and Swin-L [8] are the SOTA training-time defenses for each dataset, respectively.
-As listed in Table II, the model architectures of the SOTA training-time defenses differ from the backbone models, so our experiments are a reasonable way to evaluate transferability.
-Test-time defenses: We used DiffPure [9] as the benchmark test-time defense.
-The inputs (regardless of original or malicious) were purified by DiffPure and then fed into RaWideResNet-70-16[7] or Swin-L [8] in accordance with the dataset used.
-Preemptive defenses.
-We compared Fast Preemption with two benchmark preemptive defenses: the SOTA preemptive defense, Bi-Level [10], and the latest preemptive defense, $A^{5}$ [11].
-We used the best defense recipe, “RC (robustifier and classifier),” given by Frosio and Kautz [11].
-V-C Attacks
-Adversarial attacks.
-We evaluated defenses against two adversarial attacks: PGD [17] and AutoAttack [18].
-For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [39] and Square Attack [40]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [41]).
-Reversion attacks.
-We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective.
-Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [9]).
-V-D Evaluation metrics
-We used both clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective.
-Here, accuracy refers to the ratio of predictions that match their ground-truth labels.
-Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs.
-Additionally, to capture the image quality, the structural similarity index measure (SSIM) [42] was computed with respect to the original images.
-A larger SSIM indicates a more subtle change.
-V-E Implementation
-Our implementation partially leveraged the ART [43] and RobustBench [44] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
-VI Performance
-In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Table III.
-TABLE III: Experiment Settings.
+Table 4: Experiment Settings.
         Type         Approach         Settings
@@
                  Reversion         Attack         JPEG         Quality Factor = 50%
-3
         DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
-3
-Ours Same with our defense
+        Fast Preemption (ours)         Same with our defense
         Defense         DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
-3
-Bi-Level $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
-$\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
-3
+        Bi-Level         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
         $A^{5}$         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
-3
-Ours $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
-$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
+        Fast Preemption (ours)         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
   • 1 for CIFAR-10; 2 for ImageNet.
   • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
-TABLE IV: End-to-end performance against AutoAttack [18] and comparison with benchmark adversarial defenses.
-Dataset Defense Time Cost (s) Clean (%) Standard $L_{\infty}$ (%) Standard $L_{2}$ (%) Random $L_{\infty}$ (%)
-CIFAR-10 Training-Time RaWideResNet-70-16 [7] 0 93.3 72.8 69.3 73.4
-8
-Test-Time DiffPure [9] 0.3 91.4 75.8 83.7 76.2
-8
-Preemptive $A^{5}$ [11] 0.01 93.2 72.4 68.1 73.0
-8
-Bi-Level [10] 28 92.0 80.1 78.0 80.2
-8
-Fast Preemption (ours) 0.04 95.3 85.6 84.5 85.7
-ImageNet Training-Time Swin-L [8] 0 93.0 67.0 59.5 67.5
-8
-Test-Time DiffPure [9] 1.2 88.8 73.7 79.4 73.8
-8
-Preemptive Bi-Level [10] 43 93.6 71.5 62.8 72.3
-8
-Fast Preemption (ours) 1.3 95.1 77.1 67.4 78
-8
-Fast Preemption (ours) + DiffPure [9] 2.5 93.3 82.9 87.6 83.4
-TABLE V: Preemptive defenses resist reversion attacks.
-Original (%) Preemptive Defense Distortion Purification Preemptive Reversion (ours)
-9
-Classifier Backbone Approach Clean (%) JPEG (%) DiffPure (%) White (%) Black (%)
-93.5 Bi-Level Bi-Level Bi-Level 91.9 90.5 90.8 93.1 90.7
-9
-Bi-Level PreActResNet-18 Bi-Level 91.2 90.6 90.7 93.4 91.0
-9
-Bi-Level PreActResNet-18 Ours 91.3 90.5 90.5 93.4 90.6
-9
-ResNeXt29-32x4d PreActResNet-18 Ours 95.5 94.4 94.8 94.0 94.6
-  • Numbers in red demonstrate that only our reversion attack is feasible but only in the white-box setting.
-Figure 4: Perturbations of adversarial attacks and proposed preemptive defenses.
-Observation of differences between original image and RE show that Fast Preemption perturbations are unnoticeable.
-Comparison of perturbations between AutoAttack and Fast Preemption show that preemptive defense is more likely to succeed when perturbations are similar.
-VI-A Robustness against Adversarial Attacks
-We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [18] and victim models [7, 8].
-Table IV compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [18].
-Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive.
-For example, in comparison with the least time costly method RaWideResNet-70-16 [7] with the CIFAR-10 dataset, Fast Preemption had 2% and 12.8% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation.
-Compared with the SOTA preemptive defense, Bi-Level [10], which had the second best accuracy, Fast Preemption had 3.3% and 5.5% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster.
-Note that preemptive defenses can work together with training-time and test-time defenses.
-For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last row of Table IV.
-Figure 4 illustrates the results of qualitative analysis.
-A visual comparison of the original image and RE (average SSIM 0.962) revealed that the perturbations generated by Fast Preemption are unnoticeable.
-Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones.
-Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [45]:
-$\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
-$\delta_{Gray}=$ $\ 0.299\cdot\delta_{RGB}.R+$
-$\ 0.587\cdot\delta_{RGB}.G+0.114\cdot\delta_{RGB}.B,$
-where $\delta_{RGB}.R$, $\delta_{RGB}.G$, and $\delta_{RGB}.B$ represent the red, green, and blue channels, respectively.
-VI-B Resistance against Reversion Attacks
-We conducted reversion attacks to evaluate whether preemptive defenses are reversible.
-Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense.
-Preemptive Reversion was evaluated in both white- and black-box scenarios (see Section IV-B).
-The results are summarized in Table V.
-The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense.
-The other reversion attacks failed because they further reduced clean accuracy (due to distortion).
-Nevertheless, Preemptive Reversion is only feasible in the white-box setting.
-Without full knowledge of the defense, Fast Preemption is irreversible.
-Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails.
-However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
-Figure 5: Perturbations before and after proposed Preemptive Reversion attack. White-box reversion successfully erased most protective perturbations while black-box reversion failed.
-Moreover, the results in the fourth row of Table V, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive.
-The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion.
-However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion.
-Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
-Figure 6: Compared with the SOTA preemptive defense Bi-Level [10], our framework improved the clean and robust accuracies, and our learning algorithm reduced the time cost for defense.
-VII Ablation Studies
-We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
-The attacks were PGD [17] and AutoAttack (Standard) [18] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-As illustrated in Figure 6, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [10] in terms of clean and robust accuracies and time cost.
-These studies and results were aimed at identifying the optimal strategies.
-VII-A Classifier
-In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction.
-As demonstrated in Table VI, a classifier with higher accuracy leads to better performance.
-As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning.
-Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction.
-Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
-TABLE VI: Performance with different classifiers.
+Table 5: Performance with different classifiers.
        Ablation        Time Cost(s)        Transferable Defense
-6
        Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
@@
 The transferable defenses were all evaluated using RaWideResNet-70-16 [32] as the victim model.
-VII-B Backbone Model
+Bold indicates the best performance.
+Appendix D Experiment Settings
+The full experimental settings used for evaluations are listed in Tab. 4.
+Appendix E Implementation
+Our implementation partially leveraged the ART [30] and RobustBench [10] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
+Appendix F Ablation Studies
+We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
+The attacks were PGD [28] and AutoAttack (Standard) [9] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+As illustrated in Fig.
+6 of Sec.
+7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [29] in terms of clean and robust accuracies and time cost.
+These studies and results were aimed at identifying the optimal strategies.
+Figure 7: Different backbone models learn different features.
+PreActResNet-18 [17], WideResNet-94-16 [5], and Bi-Level [29] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
+F.1 Classifier
+In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction.
+As demonstrated in Tab.
+5, a classifier with higher accuracy leads to better performance.
+As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning.
+Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction.
+Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
+Table 6: Performance of different backbone models.
+Backbone Model: Bi-Level [29] WideResNet-94-16 [5] PreActResNet-18 [17]
+Backbone Robustness Clean (%) 89.1 93.9 88.1
+PGD (%) 43.9 77.1 61.8
+AutoAttack (%) 42.4 74.0 58.6
+Time Cost (s) for Defense 28 69 26
+White-Box Defense Clean (%) 95.5 95.8 93.8
+PGD (%) 84.9 91.6 85.9
+AutoAttack (%) 81.8 90.2 83.3
+Transferable Defense Victim: RaWideResNet-70-16 [32] Clean (%) 95.4 95.5 95.3
+PGD (%) 85.7 89.7 87.6
+AutoAttack (%) 82.7 88.0 85.9
+Transferable Defense Victim: XCiT-L12 [11] Clean (%) 95.3 95.3 95.4
+PGD (%) 72.7 78.6 79.6
+AutoAttack (%) 70.8 75.7 77.3
+  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [19] as the classifier model, and 100 steps backward propagation as Bi-Level [29].
+The transferable defenses were all evaluated using RaWideResNet-70-16 [32] and XCiT-L12 [11] as the victim models, different from the backbone models.
+Bold indicates the best performance.
+F.2 Backbone Model
 The backbone model determines the defense time cost and performance.
-As illustrated in Figure 7, different backbone models learn different features.
-To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Section I, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [13].
-Figure 7: Different backbone models learn different features.
-PreActResNet-18 [5], RaWideResNet-70-16 [7], and Bi-Level [10] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
-TABLE VII: Performance of different backbone models.
-Ablation Time Cost (s) White-box Defense Transferable Defense
-11
-Backbone Clean (%) PGD (%) AutoAttack (%) Clean (%) PGD (%) AutoAttack (%) Clean (%) PGD (%) AutoAttack (%)
-Bi-Level [10] 89.1 43.9 42.4 28 95.5 84.9 81.8 95.3 72.7 70.8
-RaWideResNet-70-16 [7] 93.5 74.5 71.4 178 95.7 91.6 89.9 95.2 78.1 76.2
-PreActResNet-18 [5] 88.1 61.8 58.6 26 93.8 85.9 83.3 95.4 79.6 77.3
-  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [37] as the classifier model, and 100 steps backward propagation as Bi-Level [10].
-The transferable defenses were all evaluated using XCiT-L12 [6] as the victim model, different from the backbone models.
-As shown in Table VII, the best adversarially trained model was RaWideResNet-70-16 [7], so it is not surprising that its white-box defense outperformed that of the other backbone models.
-However, this most robust model failed to guarantee the best transferability, and the time cost was much higher due to its complexity.
-We thus empirically selected PreActResNet-18 [5] as the backbone model in Fast Preemption as it had the least time cost and best transferability.
+As illustrated in Fig.
+7, different backbone models learn different features.
+To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec.
+1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [35].
+As shown in Tab.
+6, the best adversarially trained model was WideResNet-94-16 [5], so it is not surprising that its white-box defense outperformed that of the other backbone models.
+However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [11]), and the time cost was much higher due to its complexity.
+We thus empirically selected PreActResNet-18 [17] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [5].
 Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
-VII-C Forward-Backward (Fore-Back) Cascade Learning
-As observed in Figure 8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0).
+F.3 Forward-Backward Cascade Learning
+As observed in Fig.
+8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0).
 In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90).
 Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
-Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Figure 9.
+Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig.
+9.
 The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs.
@@
 The numbers in brackets represent robust accuracy against AutoAttack (Standard).
-VII-D Optimization Algorithm
-The results in Table VIII indicate that averaging $N$ FGSM examples results in the best performance with the least time cost.
-Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
-TABLE VIII: Performance for different $N$ and optimization algorithms.
+Table 7: Performance for different $N$ and optimization algorithms.
        Ablation        Time Cost (s)        Transferable Defense
-6
        Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
@@
        FGSM        20        0.04        95.5        87.8        85.9
-  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [7] as the victim model.
-VIII Conclusion
-We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption.
-We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths.
-In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses.
-We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised.
-This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field.
-The toolbox that will be released supports further development and offers a platform for discussing relevant technologies.
+  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [32] as the victim model. Bold indicates the best performance.
+F.4 Optimization Algorithm
+The results in Tab.
+7 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost.
+Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.

### v5

@@
 For example, in an end-to-end comparison with the state-of-the-art (SOTA) preemptive defense [32] on the ImageNet dataset [2], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [11] and was 33 times faster.
-Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [39].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible.
+Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [43].” On the basis of that assumption, Bi-Level [29] first proposed a reversion attack but failed, claiming their defense was irreversible.
+However, our reversion attack successfully defeated their defense, establishing ours as the first effective adaptive reversion attack in the white-box setting, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible.
 Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense.
@@
 Lastly, Alfarra et al.
-[2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer.
-However, the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
+[3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier.
+However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper.
+This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios.
+While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model.
+Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion).
+Furthermore, the approach of Alfarra et al.
+[3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
 Figure 2: Fast Preemption framework.
@@
 A reversion attack is used to evaluate whether a preemptive defense is reversible.
+While Bi-Level [32] first proposed a reversion attack but failed, claiming their defense was irreversible, our reversion attack successfully defeated their defense, establishing ours as the first effective reversion attack.
 In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks.
@@
 The model architectures are listed in Tab.
-3 of Appendix C.
+4 of Appendix D.
 5.2 Attacks and Defenses
@@
 The results were obtained in the settings listed in Tab.
-4 of Appendix D.
+3 of Appendix C.
 The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
@@
 8 Conclusion
-We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption.
-We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths.
-In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses.
-We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised.
-This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field.
-Acknowledgments
-This work was supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grants JPMJCR18A6 and JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3, by JST K Program Grant JPMJKP24C2 Japan, and by the project for the development and demonstration of countermeasures against disinformation and misinformation on the Internet with the Ministry of Internal Affairs and Communications of Japan.
+We introduce an efficient and transferable preemptive adversarial defense, termed Fast Preemption.
+This approach establishes a novel layer in the defense hierarchy, Proactive Defense (ours) → Training-Time Defense → Test-Time Defense, distinct from conventional test-time and training-time defenses.
+The proposed mechanism is plug-and-play, transferable, and requires no prior knowledge, empowering users to safeguard media autonomously without reliance on third-party interventions.
+Moreover, our method addresses the efficiency limitations of proactive defenses, achieving processing speeds 700 times faster than existing approaches, a critical improvement for real-world applications where substantial delays (e.g., 28 seconds per image [32]) are infeasible.
+Finally, our approach maintains or enhances clean accuracy while delivering state-of-the-art robust accuracy, overcoming the common trade-off observed in other defense methods that degrade clean accuracy.
+Nevertheless, our defense shows limitations against the proposed reversion attack in the white-box setting.
+While we design the defense allowing user customizing the backbone model (serving as a user-side key) to counter reversion attacks, simplifying user requirements is a priority for future work.
+Further research should also explore the relationship between backbone models and transferability, as discussed in Sec.
+F.2.
 Supplementary Material
-This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix C), and the experimental configurations for both attacks and defenses (Appendix D).
+This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix D), and the experimental configurations for both attacks and defenses (Appendix C).
 Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
@@
 Lastly, Alfarra et al.
-[2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer.
-However, we believe that the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
+[3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier.
+However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper.
+This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios.
+While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model.
+Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion).
+Furthermore, the approach of Alfarra et al.
+[3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
+Preemptive defenses aim to generate REs that replace original images, which are then discarded.
+These REs are more challenging to manipulate and thus resist adversarial attacks.
+Salman et al.
+[40] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples.
+They were learned using projected gradient descent (PGD) [31] and improved both in-distribution performance and robustness against unforeseen perturbations.
+Moon et al.
+[32] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy.
+However, its time cost is huge, and its transferability is low.
+Frosio and Kautz [17] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained.
+It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
+Appendix B Algorithms of Fast Preemption and
+Preemptive Reversion
+The complete defense algorithm of Fast Preemption is presented in Algorithm 1.
+With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
+The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
 Algorithm 1 Fast Preemption
@@
 5:return $\tilde{x}$
-Preemptive defenses aim to generate REs that replace original images, which are then discarded.
-These REs are more challenging to manipulate and thus resist adversarial attacks.
-Salman et al.
-[36] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples.
-They were learned using projected gradient descent (PGD) [28] and improved both in-distribution performance and robustness against unforeseen perturbations.
-Moon et al.
-[29] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy.
-However, its time cost is huge, and its transferability is low.
-Frosio and Kautz [15] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained.
-It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
-Appendix B Algorithms of Fast Preemption and
-Preemptive Reversion
-The complete defense algorithm of Fast Preemption is presented in Algorithm 1.
-With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
-The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
-Appendix C Model Architectures
+Table 3: Experiment Settings.
+       Type        Approach        Settings
+               Adversarial        Attack        PGD [31]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
+       AutoAttack [11]
+       CW [7]
+       DeCoW [28]
+       MCG [53]        Maximum Queries = 20,000
+               Reversion        Attack        JPEG        Quality Factor = 50%
+       DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
+       Fast Preemption (ours)        Same with our defense
+       Defense        DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
+       Bi-Level [32]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
+       $A^{5}$ [17]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
+       Fast Preemption (ours)        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
+  • 1 for CIFAR-10; 2 for ImageNet.
+  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
+Appendix C Experiment Settings
+The full experimental settings used for evaluations are listed in Tab. 3.
+Appendix D Model Architectures
 We evaluated attacks and defenses using models with various architectures, as listed in Tab. 4.
@@
       DeiT Base [45] (classifier)
-Table 4: Experiment Settings.
-        Type         Approach         Settings
-        Adversarial         PGD         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
-        Attack         AutoAttack
-                 Reversion         Attack         JPEG         Quality Factor = 50%
-        DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
-        Fast Preemption (ours)         Same with our defense
-        Defense         DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
-        Bi-Level         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
-        $A^{5}$         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
-        Fast Preemption (ours)         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
-  • 1 for CIFAR-10; 2 for ImageNet.
-  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
+Appendix E Implementation
+Our implementation partially leveraged the ART [33] and RobustBench [12] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
+Appendix F Ablation Studies
+We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
+The attacks were PGD [31] and AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+As illustrated in Fig.
+6 of Sec.
+7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [32] in terms of clean and robust accuracies and time cost.
+These studies and results were aimed at identifying the optimal strategies.
 Table 5: Performance with different classifiers.
@@
 Bold indicates the best performance.
-Appendix D Experiment Settings
-The full experimental settings used for evaluations are listed in Tab. 4.
-Appendix E Implementation
-Our implementation partially leveraged the ART [30] and RobustBench [10] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
-Appendix F Ablation Studies
-We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
-The attacks were PGD [28] and AutoAttack (Standard) [9] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-As illustrated in Fig.
-6 of Sec.
-7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [29] in terms of clean and robust accuracies and time cost.
-These studies and results were aimed at identifying the optimal strategies.
-Figure 7: Different backbone models learn different features.
-PreActResNet-18 [17], WideResNet-94-16 [5], and Bi-Level [29] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
 F.1 Classifier
@@
 1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [39].
+Figure 7: Different backbone models learn different features.
+PreActResNet-18 [19], WideResNet-94-16 [6], and Bi-Level [32] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
 As shown in Tab.
@@
 Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
+Appendix G Evaluation across More Models
+In the main paper, we concentrated on comparisons with top-performing adversarial training methods [6, 29].
+In this section, we extend our evaluation to include defenses tested across nine additional victim models on CIFAR-10 [26] from RobustBench [12] in a transferable context.
+These models encompass diverse architectures representative of platforms such as Twitter and Facebook, differing from both the classifier [21] and the backbone model [19].
+As shown in Tab.
+8, our method demonstrates significant performance gains compared to no proactive defense and the SOTA proactive defense [32].
+Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively.
+Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
+Table 8: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+All proactive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
+Model Metric No Proactive Defense Bi-Level [32] Fast Preemption (ours)
+Defense Learning Backbone N/A Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
+Time Cost (s) N/A 28 0.35 0.04
+RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
+Robust 71.4 81.3 86.8 85.9
+WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
+Robust 71.2 80.9 86.4 86.0
+WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
+Robust 62.1 76.2 78.7 80.8
+WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
+Robust 64.8 77.7 81.3 83.2
+PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
+Robust 42.4 54.7 55.4 61.2
+ResNest152 [41] Clean 86.9 87.6 91.2 92.4
+Robust 62.4 73.2 75.8 78.2
+ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
+Robust 52.0 68.1 66.7 72.0
+ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
+Robust 51.4 67.5 67.8 72.8
+XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
+Robust 57.9 71.3 74.6 76.0
+Average Clean 89.9 89.8 93.4 94.1
+Robust 59.5 72.3 74.8 77.3
+  • Bold indicates the best performance.
+Appendix H Defense against More Attacks
+Our primary evaluation adheres to the RobustBench [12] protocol, employing the strongest attack, AutoAttack [11].
+As detailed in Sec.
+5.2, this includes tests with standard and randomized AutoAttack variants, encompassing white-box APGD [11], FAB [10], and black-box Square [4] attacks, alongside EOT settings [5].
+In this section, we extend the evaluation to include three additional attacks: white-box Carlini-Wagner (CW) [7], black-box Deformation-Constrained Warping (DeCoW) [28], and black-box Meta Conditional Generator (MCG) [53], demonstrating the broad applicability and generalizability of our defense against diverse attack strategies.
+As shown in Tab.
+9, our defense achieves robust accuracy of 90.3%, 93.6%, and 87.3% against CW, DeCoW, and MCG attacks, respectively, outperforming SOTA adversarial training [6] and proactive defenses [32].
+Notably, these accuracies exceed those obtained against AutoAttack, further highlighting the robustness and reliability of our primary evaluation.
+Table 9: Robust accuracy (%) against various adversarial attacks with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+      Threat Model       Attack       WideResNet-94-16 [6]       Bi-Level [32]       Fast Preemption (ours)
+      White-Box       AutoAttack (Standard) [11]       74.0       82.1       86.3
+      CW [7]       80.8       85.4       90.3
+      Black-Box       DeCoW [28]       88.8       89.7       93.6
+      MCG [53]       76.1       83.3       87.3
+  • Bold indicates the best performance.
+Appendix I Visual Quality
+The noise budgets utilized (as detailed in Tab.
+3) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection.
+As illustrated in Fig.
+10, for 1,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images.
+Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
+Figure 10: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA proactive defense [32].
+All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images.
+SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
+I.1 Structural Similarity Index (SSIM)
+The SSIM is a metric for measuring similarity between two images.
+It considers luminance, contrast, and structural information.
+The SSIM between two images $x$ (reference) and $y$ (distorted) is defined as:
+$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (15)
+where:
+  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
+  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
+  • $\sigma_{xy}$: Covariance between $x$ and $y$.
+  • $C_{1}$ and $C_{2}$: Small constants added to stabilize the metric in the presence of weak denominators, defined as:
+    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
+    where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit images), and $K_{1}$ and $K_{2}$ are small positive constants (commonly, $K_{1}=0.01$ and $K_{2}=0.03$).
+The SSIM value ranges from -1 to 1, where:
+  • 1: Perfect similarity.
+  • 0: No similarity.
+  • Negative values indicate dissimilarity.
+SSIM is typically calculated locally using a sliding window across the image, and the overall SSIM is the average of local SSIM values.
+Higher SSIM values denote subtler deviations from the reference images.
+I.2 Learned Perceptual Image Patch Similarity (LPIPS)
+The LPIPS is a perceptual similarity measure that evaluates the perceptual distance between two images.
+It uses deep neural network features to compare the visual content and correlates better with human perception than traditional metrics like SSIM.
+The LPIPS distance between two images $x$ (reference) and $y$ (distorted) is computed as:
+$LPIPS(x,y)=$ (17)
+$\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}w_{l}\|\hat{\phi}_{l}(x)_{h,w}-\hat{\phi}_{l}(y)_{h,w}\|_{2}^{2},$
+where:
+  • $l$: Layer index in the deep network.
+  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
+  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
+  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
+  • $w_{l}$: Learned scalar weights for each layer $l$ (trained to match human judgments).
+LPIPS uses a pretrained deep neural network to extract feature maps from intermediate layers.
+The metric normalizes the feature maps and computes the difference between them, weighted according to the learned weights $w_{l}$.
+The LPIPS value typically ranges from 00 (perfect perceptual similarity) to higher values indicating greater dissimilarity.
+LPIPS provides a perceptually grounded comparison, making it especially useful for evaluating subtle changes in visual quality.

### v6

@@
 Abstract
-Deep learning technology has brought convenience and advanced developments but has become untrustworthy due to its sensitivity to adversarial attacks.
-Attackers may utilize this sensitivity to manipulate predictions.
-To defend against such attacks, existing anti-adversarial methods typically counteract adversarial perturbations post-attack, while we have devised a proactive strategy that preempts by safeguarding media upfront, effectively neutralizing potential adversarial effects before the third-party attacks occur.
-This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features.
-A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting.
-This strategy offers state-of-the-art transferability and protection across various systems.
-With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses.
-We have also devised the first, to our knowledge, effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised.
-This work provides a new direction to developing proactive defenses against adversarial attacks.
+Deep learning has advanced significantly but remains vulnerable to adversarial attacks, compromising its reliability.
+While conventional defenses typically mitigate perturbations post-attack, few studies explore proactive strategies for preemptive protection.
+This paper proposes Fast Preemption, a novel defense that neutralizes adversarial effects before third-party attacks.
+By employing distinct models for input labeling and feature extraction, Fast Preemption enables an efficient, transferable preemptive defense with state-of-the-art robustness across diverse systems.
+To further enhance efficiency, we introduce a forward-backward cascade learning algorithm that generates protective perturbations, leveraging forward propagation for rapid convergence and iterative backward propagation to mitigate overfitting.
+Executing in just three iterations, Fast Preemption surpasses existing training-time, test-time, and preemptive defenses.
+Additionally, we propose the first effective white-box adaptive reversion attack to evaluate the reversibility of preemptive defenses, demonstrating that our approach remains secure unless the backbone model, algorithm, and settings are entirely compromised.
+This work establishes a new paradigm for proactive adversarial defense.
 1 Introduction
-Deep learning technologies have been demonstrated to be sensitive to adversarial attacks [18].
-An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos.
-For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access the face authentication system [47].
-The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [15].
-Defending against such attacks is a critical but unsolved challenge.
-Passive defenses [6, 29, 34, 22], on the one hand, trade off classification accuracy for resistance against threats [54], but may be vulnerable to counterattacks [7, 46].
-Proactive defenses [32, 17], on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications.
-Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is untransferable to other systems.
-Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [39].
+Deep learning models are highly susceptible to adversarial attacks [18], enabling adversaries to manipulate media such as images, audio, and video.
+For instance, an attacker can alter a facial image’s identity without perceptible changes, bypassing facial authentication systems [47], or introduce imperceptible noise to traffic signs, potentially compromising autonomous driving safety [15].
+Defending against such attacks remains a critical yet unresolved challenge.
+Passive defenses [6, 29, 34, 22] enhance robustness at the cost of classification accuracy [54] and remain susceptible to adaptive attacks [7, 46].
+Proactive defenses [32, 17], while mitigating threats preemptively, often lack real-time efficiency and require paired protection-inference modules, limiting transferability.
+Moreover, users relying on third-party models must contend with potential backdoor threats [39].
 Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
-To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Fig.
+To address these challenges, we propose a proactive defense strategy, Fast Preemption, as illustrated in Fig.
 1.
-Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party.
-Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection.
-In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized.
-This strategy has three distinct advantages.
-(i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system.
-(ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system [16, 39].
-(iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated within a classification model) is breached, all inputs are affected.
-In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings.
-Therefore, Fast Preemption meets four basic requirements:
+Fast Preemption preemptively “attacks” media on the user side to enhance robustness before potential third-party attacks.
+Specifically, it learns robust features, predicts possible adversarial perturbations, and then amplifies these features while reversing the predicted perturbations for protection.
+As a result, any subsequent adversarial attack is inherently neutralized.
+This approach offers three key advantages.
+(i) Unlike passive defenses, Fast Preemption enhances threat resistance without compromising classification accuracy and requires no modifications to the classification system.
+(ii) It supports plug-and-play deployment without paired protection-inference modules, mitigating security and privacy concerns associated with external model providers [16, 39].
+(iii) Traditional training- and test-time defenses, integrated within classification models, risk full-system compromise if breached.
+In contrast, Fast Preemption localizes vulnerabilities to individual users, as each employs distinct configurations.
+Thus, Fast Preemption satisfies four fundamental requirements:
   • Preserve clean accuracy and improve robust accuracy: Basic requirements for all defenses.
@@
   • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
-We propose the Fast Preemption defense includes three components: a classifier model for automatically labeling inputs, a backbone model for learning features (differ from the classifier), and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations.
-We conducted evaluation on two datasets [26, 2] across six models [19, 6, 35, 29, 32, 13], defending against two typical adversarial attacks [31, 11].
-The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [6, 29, 34, 32, 17].
-For example, in an end-to-end comparison with the state-of-the-art (SOTA) preemptive defense [32] on the ImageNet dataset [2], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [11] and was 33 times faster.
-Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [43].” On the basis of that assumption, Bi-Level [29] first proposed a reversion attack but failed, claiming their defense was irreversible.
-However, our reversion attack successfully defeated their defense, establishing ours as the first effective adaptive reversion attack in the white-box setting, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible.
-Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense.
-Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier).
-If clean accuracy increases after reversion, the reversion succeeded.
-If it decreases after reversion (due to distortion), reversion failed.
-The results from reversion attack testing demonstrate that the Preemptive Reversion attack can smartly erase protective perturbations when all defense details are compromised (i.e., a white-box situation), whereas blind JPEG distortion and adversarial purification [34] only further distort an image.
-However, this white-box situation is impractical with the Fast Preemption strategy for real-world applications because the proposed framework enables users to individually customize their defense settings.
-Apart from a white-box situation, Fast Preemption protection cannot be reversed using existing reversion approaches.
-We summarize our contributions as follows: (i) We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module.
-It is the first to use different models for the classifier and backbone.
-(ii) We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction.
-(iii) We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
-2 Related Work
-Existing adversarial defenses can be categorized into three primary types: training-time, test-time, and preemptive defenses. Appendix A provides a more detailed review.
-Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels [18, 42, 41, 19, 13, 50, 14, 35, 54, 6, 29, 44].
-However, they have two shortcomings.
-First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness.
-Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy.
-Test-time defenses are divided to detection and purification.
-Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs [9, 56, 27, 22].
-However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks.
-Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
-In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations.
-Earlier studies used image distortion for purification [48, 20, 37].
-Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious [38, 34].
-The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [7, 46].
-Preemptive defenses seek to generate robust examples (REs) that replace the original images, which are subsequently discarded [40, 32, 17].
-These REs are inherently more difficult to manipulate, thereby providing resistance against adversarial attacks.
-However, such approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, lower robustness compared to other defense mechanisms.
-Lastly, Alfarra et al.
-[3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier.
-However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper.
-This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios.
-While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model.
-Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion).
-Furthermore, the approach of Alfarra et al.
-[3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
+We introduce Fast Preemption, a preemptive defense comprising three key components: a classifier model for automatic input labeling, a backbone model (distinct from the classifier) for feature learning, and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations.
+We evaluate Fast Preemption on two datasets and thirteen models, defending against five representative adversarial attacks.
+Results demonstrate its superiority in clean and robust accuracy, transferability, and computational efficiency over benchmark training-time, test-time, and preemptive defenses [6, 29, 34, 32, 17].
+For example, on CIFAR-10 [26], Fast Preemption outperformed the state-of-the-art (SOTA) preemptive defense [32], achieving 2.5% and 5.1% higher clean and robust accuracy, respectively, against AutoAttack (Standard) [11], while being 700 times faster.
+Shannon famously stated, “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them” [43].
+Following this principle, Bi-Level [32] first introduced a reversion attack but failed, asserting their defense was irreversible.
+In contrast, our reversion attack successfully circumvented their defense, making ours the first effective adaptive reversion attack in the white-box setting, termed Preemptive Reversion, to assess the reversibility of preemptive defenses.
+Preemptive Reversion applies a secondary preemptive defense to the already protected example and then reverses the newly introduced perturbations.
+Since the secondary defense generates perturbations similar to the first under identical conditions, this process effectively tests a defense’s resilience to reversion.
+Additionally, we propose a more rigorous evaluation protocol, which mandates a decrease in clean accuracy after preemptive defense (achieved by selecting a suboptimal classifier).
+If clean accuracy improves post-reversion, the attack is successful; if it further deteriorates due to distortion, reversion has failed.
+Results of reversion attacks show that Preemptive Reversion effectively removes protective perturbations in a white-box setting, where all defense details are exposed.
+In contrast, blind JPEG distortion and adversarial purification [34] further degrade the image rather than neutralizing protections.
+However, this white-box assumption is unrealistic for real-world applications of Fast Preemption, as the framework allows users to customize their defense settings individually.
+Beyond the white-box scenario, existing reversion methods fail to circumvent Fast Preemption’s protection.
+We summarize our contributions as follows: (i) We introduce Fast Preemption, a novel preemptive defense that enhances transferability without requiring ground truths or paired modules.
+It is the first approach to utilize distinct models for the classifier and backbone.
+(ii) We propose the forward-backward (Fore-Back) cascade learning algorithm, which accelerates preemptive defense, enabling efficient user interaction.
+(iii) We develop Preemptive Reversion, the first effective adaptive reversion attack for assessing the robustness of preemptive defenses in a white-box setting.
+Additionally, we establish a more comprehensive evaluation protocol to determine reversion success.
+2 Related Works
+Existing adversarial defenses can be classified into three main categories: training-time, test-time, and preemptive defenses. A detailed review is provided in Appendix F.
+Training-time defenses aim to develop robust deep learning models that can correctly classify adversarial examples (AEs) based on their ground truth labels [18, 42, 41, 19, 13, 50, 14, 35, 54, 6, 29, 44].
+However, they have two key limitations.
+First, they require large amounts of high-quality, diverse data for training.
+Second, training-time defenses often sacrifice clean accuracy to improve robust accuracy.
+Test-time defenses are typically divided into detection and purification.
+Detection models identify AEs by distinguishing them from benign inputs [9, 56, 27, 22], but detection is challenging for unseen attacks, as detectors are often trained on known AEs.
+Furthermore, detection may misclassify benign images as AEs, reducing clean accuracy.
+In contrast, adversarial purification aims to remove perturbations in AEs.
+Early approaches used image distortion for purification [48, 20, 37], while recent methods generate benign replacements regardless of input type [38, 34].
+A key limitation of purification is that the difference between original and generated examples can degrade clean accuracy, and counterattack potential remains [7, 46].
 Figure 2: Fast Preemption framework.
@@
 The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
+Preemptive defenses aim to generate robust examples (REs) that replace the original images, which are then discarded [40, 32, 17].
+These REs are inherently more resilient to manipulation, offering resistance against adversarial attacks.
+However, these approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, reduced robustness compared to other defense mechanisms.
+Alfarra et al.
+[3] proposed a test-time defense incorporating preemptive protection as an anti-adversary layer guided by a standard classifier.
+However, their method is inherently vulnerable if the classifier is compromised, a limitation they explicitly acknowledge.
+This weakness undermines its practical viability in test-time scenarios.
+While they describe this setting as the “least realistic,” we argue that it aligns with the standard white-box attack model, where adversaries have full access to the classifier.
+Under even stricter assumptions, adversaries may also be aware of the defense mechanism itself (white-box reversion).
+Moreover, their approach relies solely on forward propagation learning, increasing susceptibility to overfitting.
 3 Preemptive Defense: Fast Preemption
@@
 In summary, the forward and backward propagation both enhance robustness, but in different optimization directions.
-The complete algorithm is presented in Algorithm 1 of Appendix B.
+The complete algorithm is presented in Algorithm 1 of Appendix D.
 3.3.1 Loss Function
@@
 8 is AE $x_{t}^{a}$.
-Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
+Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next RE, $x_{t+1}$:
 $x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}-\eta\cdot(x_{t}^{a}-x_{t}))$ (9)
 $=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))).$
-The SOTA preemptive defense strategy [32] conducts only backward propagation learning.
-However, adversarial attacks can be randomly initialized, so backward propagation requires many steps to converge (minimum ten steps).
-The Fast Preemption strategy accelerates defense by running forward propagation before running backward propagation learning.
+The SOTA preemptive defense strategy [32] relies solely on backward propagation learning, which requires multiple steps to converge due to the random initialization of adversarial attacks (at least ten steps).
+In contrast, the Fast Preemption strategy accelerates defense by incorporating forward propagation before backward propagation learning.
 3.3.3 Combining Forward and Backward Propagation
@@
 $s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
-the final output of which is robust example $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
+the final output of which is RE $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
 $x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}+\eta\cdot(x_{t}^{r}-x_{t}))$ (11)
@@
 4 Adaptive Attack: Preemptive Reversion
-A reversion attack is used to evaluate whether a preemptive defense is reversible.
-While Bi-Level [32] first proposed a reversion attack but failed, claiming their defense was irreversible, our reversion attack successfully defeated their defense, establishing ours as the first effective reversion attack.
+A reversion attack is used to assess the reversibility of a preemptive defense.
+While Bi-Level [32] first proposed a reversion attack but failed to reverse their defense, claiming it was irreversible, our reversion attack successfully overcame their defense, making ours the first effective reversion attack.
 In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks.
-Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario.
-Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
+Preemptive Reversion is the only viable approach to reverse preemptive defenses in the white-box scenario.
+However, without full knowledge of the backbone model and settings, our Fast Preemption defense remains difficult to reverse, ensuring its reliability.
+Figure 3: Illustration of Preemptive Reversion algorithm.
+It first conducts a secondary preemptive defense on RE $x^{r}$ and then reverses the perturbations of the secondary defense: $(x^{r})^{r}-x^{r}$.
 4.1 Algorithm
@@
 3.
-The complete algorithm is presented in Algorithm 2 of Appendix B.
+The complete algorithm is presented in Algorithm 2 of Appendix D.
 Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$.
@@
 $\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
-Figure 3: Illustration of Preemptive Reversion algorithm.
-It first conducts a secondary preemptive defense on RE $x^{r}$ and then reverses the perturbations of the secondary defense: $(x^{r})^{r}-x^{r}$.
 4.2 Attack Scenarios
-We define two scenarios for a reversion attack in accordance with the knowledge the attacker has about the backbone model and defense algorithm:
-  • White-Box: The attacker knows both the backbone model and defense algorithm (including settings), so the attacker can conduct a secondary defense the same as the user does.
-This is the strongest, yet impractical, assumption because different users may use different settings.
-  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (infinite selections for users).
-The challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
-Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our Fast Preemption defense, but it is feasible only in the white-box scenario.
-Without full knowledge of the defense, Fast Preemption is reliable.
+We define two reversion attack scenarios via the attacker’s knowledge of the backbone model and defense algorithm:
+  • White-Box: The attacker has complete knowledge of both the backbone model and defense algorithm (including settings), enabling them to replicate the secondary defense.
+This scenario, while the strongest, is impractical due to varying user settings.
+  • Black-Box: The attacker knows the defense algorithm (typically public) but lacks knowledge of the backbone model, which can vary across users.
+The challenge in this scenario is that the features learned in the secondary defense may differ from those learned by the user.
+Our proposed Preemptive Reversion attack is the first effective method for evaluating the reversibility of preemptive defenses, including Fast Preemption, but it is only feasible in the white-box scenario.
+Without full knowledge of the defense, Fast Preemption remains reliable.
 4.3 Evaluation Protocol
-If a reversion attack succeeds, the improved clean accuracy should drop.
-However, image distortion can create a similar effect.
-Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack.
-To address this issue, we proposed a comprehensive protocol that can distinguish successful reversion from distortion.
-First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy.
-In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded.
-Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
-5 Experiments
-5.1 Datasets, Classifier, and Backbone Models
+If a reversion attack succeeds, clean accuracy should decrease.
+However, image distortion can lead to a similar outcome, meaning a drop in clean accuracy does not necessarily indicate a successful reversion.
+To address this, we propose a comprehensive protocol to distinguish successful reversion from distortion.
+First, we intentionally assign incorrect labels to partial inputs, ensuring the clean accuracy after defense is lower than the original.
+If clean accuracy approaches the original value, the reversion attack is successful; otherwise, the drop in accuracy is due to distortion.
+5 Experiment Settings
+Datasets, Classifier, and Backbone Models.
 We conducted experiments on two datasets: CIFAR-10 [26] and ImageNet [2].
@@
 The model architectures are listed in Tab.
-4 of Appendix D.
-5.2 Attacks and Defenses
+9 of Appendix E.
 Adversarial attacks.
-We evaluated defenses against two adversarial attacks: PGD [31] and AutoAttack [11].
-For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [10] and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]).
+We evaluated defenses against five adversarial attacks: white-box PGD [31], AutoAttack [11], and Carlini-Wagner (CW) attack [7]; black-box Deformation-Constrained Warping (DeCoW) [28] and Meta Conditional Generator (MCG) [53].
+For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, FAB [10], and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]).
 Reversion attacks.
-We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective.
-Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [34]).
-Training-time defenses: We compared Fast Preemption with WideResNet-94-16 [6] and Swin-L [29] on the CIFAR-10 and ImageNet datasets, respectively.
-The model architectures of these defenses differ from the backbone models.
-Test-time defenses: We used DiffPure [34] as the benchmark test-time defense.
-The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [29] in accordance with the dataset used.
-Preemptive defenses.
-We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [32], and the latest preemptive defense, $A^{5}$ [17].
+We examined: (i) whether protected samples induced by preemptive defenses cannot be easily reversed to their original unprotected state, which we define as “irreversible”; and (ii) the effectiveness of the proposed adaptive reversion attack, Preemptive Reversion.
+Along with Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [34]).
+Training-time defenses.
+We compared Fast Preemption with WideResNet-94-16 [6] and Swin-L [29] on the CIFAR-10 and ImageNet datasets, respectively.
+The architectures of these models differ from the backbone models.
+To assess the transferability of preemptive defenses, we conducted experiments across additional nine victim models, with results presented in Tab.
+7 of Appendix C.
+Test-time defenses. We used DiffPure [34] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [29] depending on the dataset.
+Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [32], and the latest preemptive defense, $A^{5}$ [17].
 5.1 Evaluation metrics
-We used clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective.
-Here, accuracy refers to the ratio of predictions that match their ground-truth labels.
-Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs.
+We evaluated performance using clean accuracy and robust accuracy, where higher values indicate better defense effectiveness.
+Clean accuracy was computed for benign images (original and REs), while robust accuracy was computed for AEs.
+Additionally, to assess image quality, we calculated the structural similarity index measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] relative to the original images.
+Higher SSIM and lower LPIPS values indicate subtler changes in the images.
 Table 1: End-to-end performance against AutoAttack [11] and comparison with benchmark adversarial defenses.
@@
   • Bold indicates the best performance.
+Table 2: Robust accuracy (%) against various adversarial attacks with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+Threat Model Attack WideResNet-94-16 Bi-Level Ours
+White-Box AutoAttack 74.0 82.1 86.5
+CW 80.8 85.4 90.3
+Black-Box DeCoW 88.8 89.7 93.6
+MCG 76.1 83.3 87.3
+  • AutoAttack is the standard version. Bold indicates the best performance.
 6 Performance
-In this section, we present the end-to-end evaluation results and comparisons with benchmarks.
-The results were obtained in the settings listed in Tab.
-3 of Appendix C.
-The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
+The experiments were conducted using the settings outlined in Tab. 8 of Appendix E. Time costs were recorded on an NVIDIA A100 40GB GPU.
 Table 3: Preemptive defenses resist reversion attacks.
@@
   • Numbers in red demonstrate that only our reversion attack is feasible but only in the white-box setting.
-Figure 4: Perturbations of adversarial attacks and proposed preemptive defenses.
-Observation of differences between original image and RE show that Fast Preemption perturbations are unnoticeable.
-Comparison of perturbations between AutoAttack and Fast Preemption show that preemptive defense is more likely to succeed when perturbations are similar.
+Figure 4: Perturbations of adversarial attacks and the proposed preemptive defenses.
+Visual comparison between the original image and the RE demonstrates that Fast Preemption perturbations are imperceptible.
+Additionally, comparison of perturbations between AutoAttack and Fast Preemption reveals that preemptive defense is more likely to succeed when perturbations are similar.
 6.1 Robustness against Adversarial Attacks
-We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [11] and victim models [6, 29].
-Table 1 compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [11].
-Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive.
-For example, in comparison with the least time costly method WideResNet-94-16 [6] with the CIFAR-10 dataset, Fast Preemption had 1.6% and 11.9% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation.
-Compared with the SOTA preemptive defense, Bi-Level [32], which had the second best accuracy, Fast Preemption had 2.5% and 5.1% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster.
-Note that preemptive defenses can work together with training-time and test-time defenses.
-For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last rows of CIFAR-10 and ImageNet in Tab.
-1, respectively.
-Figure 4 illustrates the results of qualitative analysis.
-A visual comparison of the original image and RE revealed that the perturbations generated by Fast Preemption are unnoticeable.
-Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones.
-Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [30]:
+We evaluated transferable defense performance under the condition where the defense lacks knowledge of both the attacks and victim models [6, 29].
+Table 1 and Tab.
+2 compares Fast Preemption with benchmark defenses against four attacks.
+Fast Preemption outperformed the benchmarks in both clean and robust accuracies while maintaining competitive time costs.
+For instance, compared to the least time-costly method, WideResNet-94-16 [6] on CIFAR-10 under Standard $L_{\infty}$ AutoAttack, Fast Preemption achieved 1.6% and 11.9% higher clean and robust accuracies, respectively, with only a marginal increase of 0.04 seconds per generation.
+Compared to the SOTA preemptive defense Bi-Level [32], which had the second-best accuracy, Fast Preemption improved clean and robust accuracies by 2.5% and 5.1%, respectively, while being 700 times faster.
+Notably, preemptive defenses can be combined with training- and test-time defenses.
+For example, when paired with a test-time defense, Fast Preemption’s performance was significantly enhanced, as shown in the final rows of CIFAR-10 and ImageNet in Tab.
+1.
+Figure 4 provides a qualitative analysis, comparing the original images with their corresponding REs.
+The perturbations introduced by Fast Preemption are barely perceptible.
+The average SSIM and LPIPS scores for 2,000 examples from CIFAR-10 and ImageNet are 0.97 and 0.01, respectively, both surpassing those of standard AutoAttack and the SOTA preemptive defense Bi-Level [32] (more details in Appendix A).
+Additionally, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations closely resemble those of the attack.
+The perturbation illustration $\delta_{Gray}$ was generated by normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$, converting $\delta_{RGB}$ to grayscale using the NTSC formula [30].
 $\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
@@
 6.2 Resistance against Reversion Attacks
-We conducted reversion attacks to evaluate whether preemptive defenses are reversible.
-Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense.
-Preemptive Reversion was evaluated in both white- and black-box scenarios (see Sec.
+We conducted reversion attacks to assess the reversibility of preemptive defenses.
+In addition to the proposed Preemptive Reversion attack, we employed JPEG compression and adversarial purification, which do not require knowledge of the defense.
+Preemptive Reversion was tested in both white-box and black-box scenarios (see Sec.
 4.2).
-The results are summarized in Tab.
-2.
-The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense.
-The other reversion attacks failed because they further reduced clean accuracy (due to distortion).
-Nevertheless, Preemptive Reversion is only feasible in the white-box setting.
-Without full knowledge of the defense, Fast Preemption is irreversible.
-Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails.
-However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
+The results, summarized in Tab.
+3, reveal that only Preemptive Reversion, indicated in red, succeeded in reversing the preemptive defense, bringing the accuracy closer to the original.
+Other reversion attacks failed, further degrading clean accuracy due to distortion.
+Notably, Preemptive Reversion is only effective in the white-box setting, as Fast Preemption remains irreversible without full knowledge of the defense.
+Figure 5 corroborates our findings, showing that the white-box reversion successfully eliminates most protective perturbations, whereas the black-box reversion fails.
+However, due to the randomization inherent in Fast Preemption (i.e., randomly initialized FGSM), the white-box reversion cannot entirely remove all protective perturbations.
+The fourth row in Tab.
+3, which was obtained without our evaluation protocol, further highlights the comprehensiveness of our evaluation method.
+The other results demonstrate that Preemptive Reversion in the white-box setting is the only reversion attack that improves clean accuracy after reversion.
+The lower numbers in the fourth row suggest that the accuracy decrease observed with Preemptive Reversion in the white-box setting resulted from successful reversion, while other accuracy declines were due to distortion.
+Consequently, a decrease in accuracy, when solely considered from the fourth row, does not necessarily indicate the success of a reversion attack.
 Figure 5: Perturbations before and after proposed Preemptive Reversion attack. White-box reversion successfully erased most protective perturbations while black-box reversion failed.
-Moreover, the results in the fourth row of Tab.
-2, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive.
-The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion.
-However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion.
-Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
 7 Ablation Studies
-We conducted ablation studies on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset.
-We used AutoAttack (Standard) [11] against a black-box target model RaWideResNet-70-16 [35], employing $L_{\infty}$-norm perturbations with $\epsilon=8/255$.
-As demonstrated in Fig.
-6, Fast Preemption significantly outperforms the SOTA Bi-Level method [32], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to just 0.04 seconds.
-These studies were conducted to identify the optimal configurations, with more extensive ablation results presented in Appendix F.
+We conducted an ablation study on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset and AutoAttack (Standard) [11] ($L_{\infty}$-norm $\epsilon=8/255$) against a black-box target model, RaWideResNet-70-16 [35].
+As shown in Fig.
+6, Fast Preemption significantly outperforms the SOTA Bi-Level method [32], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to 0.04 seconds.
+These studies were conducted to identify optimal configurations, with further ablation results provided in Appendix B.
 Figure 6: Compared with the SOTA preemptive defense Bi-Level [32], our framework improved the clean and robust accuracies, and our learning algorithm reduced the time cost for defense.
 8 Conclusion
-We introduce an efficient and transferable preemptive adversarial defense, termed Fast Preemption.
-This approach establishes a novel layer in the defense hierarchy, Proactive Defense (ours) → Training-Time Defense → Test-Time Defense, distinct from conventional test-time and training-time defenses.
-The proposed mechanism is plug-and-play, transferable, and requires no prior knowledge, empowering users to safeguard media autonomously without reliance on third-party interventions.
-Moreover, our method addresses the efficiency limitations of proactive defenses, achieving processing speeds 700 times faster than existing approaches, a critical improvement for real-world applications where substantial delays (e.g., 28 seconds per image [32]) are infeasible.
-Finally, our approach maintains or enhances clean accuracy while delivering state-of-the-art robust accuracy, overcoming the common trade-off observed in other defense methods that degrade clean accuracy.
-Nevertheless, our defense shows limitations against the proposed reversion attack in the white-box setting.
-While we design the defense allowing user customizing the backbone model (serving as a user-side key) to counter reversion attacks, simplifying user requirements is a priority for future work.
-Further research should also explore the relationship between backbone models and transferability, as discussed in Sec.
-F.2.
+We propose Fast Preemption, an efficient and transferable preemptive adversarial defense that introduces a new layer in the defense hierarchy: Proactive Defense (ours) → Training-Time Defense → Test-Time Defense.
+Unlike conventional training- and test-time defenses, Fast Preemption is plug-and-play, transferable, and requires no prior knowledge, enabling users to safeguard media independently without relying on third-party interventions.
+Our approach also addresses the efficiency limitations of existing proactive defenses, achieving processing speeds 700 times faster than prior methods, an essential improvement for real-world applications, where significant delays (e.g., 28 seconds per image [32]) are impractical.
+Moreover, Fast Preemption maintains or improves clean accuracy while achieving SOTA robust accuracy, overcoming the common trade-off where defenses degrade clean accuracy.
+Despite its advantages, our defense exhibits limitations against the proposed reversion attack in the white-box setting.
+While allowing users to customize the backbone model as a user-side key mitigates reversion attacks, further work is needed to simplify user requirements.
+Future research should also investigate the relationship between backbone models and transferability, as discussed in Sec.
+B.2.
 Supplementary Material
-This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix D), and the experimental configurations for both attacks and defenses (Appendix C).
-Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
-Appendix A Related Work
+This supplementary material provides a comprehensive analysis of visual quality (Appendix A) and presents an extensive ablation study (Appendix B), investigating critical components such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
+Additionally, it includes evaluation results against a broader set of models (Appendix C), elaborates on the algorithms underlying the proposed adversarial defense and adaptive reversion attack (Appendix D), and details the experimental configurations for both attack and defense scenarios (Appendix E).
+Appendix A Visual Quality
+The noise budgets utilized (as detailed in Tab.
+8) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection.
+As illustrated in Fig.
+7, for 2,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images.
+Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
+Figure 7: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA proactive defense [32].
+All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images.
+SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
+A.1 Structural Similarity Index (SSIM)
+The SSIM is a metric for measuring similarity between two images.
+It considers luminance, contrast, and structural information.
+The SSIM between two images $x$ (reference) and $y$ (distorted) is defined as:
+$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (15)
+where:
+  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
+  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
+  • $\sigma_{xy}$: Covariance between $x$ and $y$.
+  • $C_{1}$ and $C_{2}$: Small constants added to stabilize the metric in the presence of weak denominators, defined as:
+    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
+    where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit images), and $K_{1}$ and $K_{2}$ are small positive constants (commonly, $K_{1}=0.01$ and $K_{2}=0.03$).
+The SSIM value ranges from -1 to 1, where:
+  • 1: Perfect similarity.
+  • 0: No similarity.
+  • Negative values indicate dissimilarity.
+SSIM is typically calculated locally using a sliding window across the image, and the overall SSIM is the average of local SSIM values.
+Higher SSIM values denote subtler deviations from the reference images.
+A.2 Learned Perceptual Image Patch Similarity (LPIPS)
+The LPIPS is a perceptual similarity measure that evaluates the perceptual distance between two images.
+It uses deep neural network features to compare the visual content and correlates better with human perception than traditional metrics like SSIM.
+The LPIPS distance between two images $x$ (reference) and $y$ (distorted) is computed as:
+$LPIPS(x,y)=$ (17)
+$\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}w_{l}\|\hat{\phi}_{l}(x)_{h,w}-\hat{\phi}_{l}(y)_{h,w}\|_{2}^{2},$
+where:
+  • $l$: Layer index in the deep network.
+  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
+  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
+  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
+  • $w_{l}$: Learned scalar weights for each layer $l$ (trained to match human judgments).
+LPIPS uses a pretrained deep neural network to extract feature maps from intermediate layers.
+The metric normalizes the feature maps and computes the difference between them, weighted according to the learned weights $w_{l}$.
+The LPIPS value typically ranges from 00 (perfect perceptual similarity) to higher values indicating greater dissimilarity.
+LPIPS provides a perceptually grounded comparison, making it especially useful for evaluating subtle changes in visual quality.
+Appendix B Ablation Studies
+We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
+The attacks were PGD [31] and AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+As illustrated in Fig.
+6 of Sec.
+7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [32] in terms of clean and robust accuracies and time cost.
+These studies and results were aimed at identifying the optimal strategies.
+Table 4: Performance with different classifiers.
+       Ablation        Time Cost(s)        Transferable Defense
+       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
+       Ground Truths        100        28        97.2        86.9        83.1
+       ResNeXt        95.9        28        95.4        85.7        82.7
+       Bi-Level        89.1        29        91.9        83.9        81.0
+  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [32].
+The transferable defenses were all evaluated using RaWideResNet-70-16 [35] as the victim model.
+Bold indicates the best performance.
+B.1 Classifier
+In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction.
+As demonstrated in Tab.
+4, a classifier with higher accuracy leads to better performance.
+As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning.
+Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction.
+Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
+Table 5: Performance of different backbone models.
+Backbone Model: Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
+Backbone Robustness Clean (%) 89.1 93.9 88.1
+PGD (%) 43.9 77.1 61.8
+AutoAttack (%) 42.4 74.0 58.6
+Time Cost (s) for Defense 28 69 26
+White-Box Defense Clean (%) 95.5 95.8 93.8
+PGD (%) 84.9 91.6 85.9
+AutoAttack (%) 81.8 90.2 83.3
+Transferable Defense Victim: RaWideResNet-70-16 [35] Clean (%) 95.4 95.5 95.3
+PGD (%) 85.7 89.7 87.6
+AutoAttack (%) 82.7 88.0 85.9
+Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
+PGD (%) 72.7 78.6 79.6
+AutoAttack (%) 70.8 75.7 77.3
+  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [21] as the classifier model, and 100 steps backward propagation as Bi-Level [32].
+The transferable defenses were all evaluated using RaWideResNet-70-16 [35] and XCiT-L12 [13] as the victim models, different from the backbone models.
+Bold indicates the best performance.
+B.2 Backbone Model
+The backbone model determines the defense time cost and performance.
+As illustrated in Fig.
+8, different backbone models learn different features.
+To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec.
+1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [39].
+Figure 8: Different backbone models learn different features.
+PreActResNet-18 [19], WideResNet-94-16 [6], and Bi-Level [32] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
+As shown in Tab.
+5, the best adversarially trained model was WideResNet-94-16 [6], so it is not surprising that its white-box defense outperformed that of the other backbone models.
+However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [13]), and the time cost was much higher due to its complexity.
+We thus empirically selected PreActResNet-18 [19] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [6].
+Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
+B.3 Forward-Backward Cascade Learning
+As observed in Fig.
+9, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0).
+In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90).
+Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
+Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig.
+10.
+The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs.
+F0-B3).
+Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
+Figure 9: Forward propagation (F) converges faster while backward propagation (B) alleviates overfitting.
+The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
+The numbers in brackets represent robust accuracy against AutoAttack (Standard).
+“Alternate” represents the running of forward and backward propagation alternatively for 100 steps in total.
+Figure 10: By combining forward (F) and backward (B) propagation, Fast Preemption obtains performance comparable to that of the SOTA preemptive defense [32] by using only three steps (F1–B2).
+The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
+The numbers in brackets represent robust accuracy against AutoAttack (Standard).
+Table 6: Performance for different $N$ and optimization algorithms.
+       Ablation        Time Cost (s)        Transferable Defense
+       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
+       PGD        1        0.7        95.3        87.5        85.8
+       FGSM        1        0.04        95.4        86.8        85.4
+       PGD        20        0.8        95.2        87.7        85.9
+       FGSM        20        0.04        95.5        87.8        85.9
+  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [35] as the victim model. Bold indicates the best performance.
+B.4 Optimization Algorithm
+The results in Tab.
+6 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost.
+Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
+Table 7: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+All proactive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
+Model Metric No Proactive Defense Bi-Level [32] Fast Preemption (ours)
+Defense Learning Backbone N/A Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
+Time Cost (s) N/A 28 0.35 0.04
+RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
+Robust 71.4 81.3 86.8 85.9
+WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
+Robust 71.2 80.9 86.4 86.0
+WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
+Robust 62.1 76.2 78.7 80.8
+WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
+Robust 64.8 77.7 81.3 83.2
+PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
+Robust 42.4 54.7 55.4 61.2
+ResNest152 [41] Clean 86.9 87.6 91.2 92.4
+Robust 62.4 73.2 75.8 78.2
+ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
+Robust 52.0 68.1 66.7 72.0
+ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
+Robust 51.4 67.5 67.8 72.8
+XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
+Robust 57.9 71.3 74.6 76.0
+Average Clean 89.9 89.8 93.4 94.1
+Robust 59.5 72.3 74.8 77.3
+  • Bold indicates the best performance.
+Appendix C Evaluation across More Models
+In the main paper, we concentrated on comparisons with top-performing adversarial training methods [6, 29].
+In this section, we extend our evaluation to include defenses tested across nine additional victim models on CIFAR-10 [26] from RobustBench [12] in a transferable context.
+These models encompass diverse architectures representative of platforms such as Twitter and Facebook, differing from both the classifier [21] and the backbone model [19].
+As shown in Tab.
+7, our method demonstrates significant performance gains compared to no proactive defense and the SOTA proactive defense [32].
+Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively.
+Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
+Appendix D Detailed Algorithms
+The complete defense algorithm of Fast Preemption is presented in Algorithm 1.
+With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
+The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
+Algorithm 1 Fast Preemption
+1:Original image $x$, classifier $f_{c}$, backbone model $f_{b}$, maximum $L_{\infty}$-norm distance $\epsilon$, learning rate $\eta$, maximum forward steps $T_{\rightarrow}$ and backward steps $T_{\leftarrow}$, number of examples for averaging perturbations $N$.
+2:Robust example $x^{r}$.
+3:// Forward Propagation
+4:$x_{0}=x$
+5:for $t<T_{\rightarrow}$ do
+6:       for $n=1,...,N$ do
+7:             $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
+8:             $x_{t,n}=x_{t}+\delta_{n}$
+9:             $x_{t,n}^{r}=\prod_{x,\epsilon}(x_{t,n}-\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM decreases $\ell$
+10:       end for
+11:       $\delta_{t}^{r}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{r}-x_{t}$$\triangleright$ averages perturbations
+12:       $x_{t+1}=x_{t}+\eta\cdot\delta_{t}^{r}$$\triangleright$ direct enhancement
+13:end for
+14:// Backward Propagation
+15:$x_{0}=x_{T_{\rightarrow}}$
+16:for $t<T_{\leftarrow}$ do
+17:       for $n=1,...,N$ do
+18:             $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
+19:             $x_{t,n}=x_{t}+\delta_{n}$
+20:             $x_{t,n}^{a}=\prod_{x,\epsilon}(x_{t,n}+\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM increases $\ell$
+21:       end for
+22:       $\delta_{t}^{a}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{a}-x_{t}$$\triangleright$ averages perturbations
+23:       $x_{t+1}=x_{t}-\eta\cdot\delta_{t}^{a}$$\triangleright$ reverses the attack
+24:end for
+25:return $x^{r}=x_{T_{\leftarrow}}$
+Algorithm 2 Preemptive Reversion
+1:Robust example $x^{r}$.
+2:Reversed example $\tilde{x}$.
+3:$(x^{r})^{r}=PreemptiveDefense(x^{r})$$\triangleright$ secondary defense
+4:$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r})$$\triangleright$ reverses the secondary defense
+5:return $\tilde{x}$
+Appendix E Experiment Settings
+The full experimental settings used for evaluations are listed in Tab.
+8.
+We evaluated attacks and defenses using models with various architectures, as listed in Tab.
+9.
+Our implementation partially leveraged the ART [33] and RobustBench [12] leaderboards by using PyTorch.
+The experimental machine was equipped with an NVIDIA A100 40GB GPU.
+Table 8: Experiment Settings.
+       Type        Approach        Settings
+               Adversarial        Attack        PGD [31]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
+       AutoAttack [11]
+       CW [7]
+       DeCoW [28]
+       MCG [53]        Maximum Queries = 20,000
+               Reversion        Attack        JPEG        Quality Factor = 50%
+       DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
+       Fast Preemption (ours)        Same with our defense
+       Defense        DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
+       Bi-Level [32]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
+       $A^{5}$ [17]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
+       Fast Preemption (ours)        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
+  • 1 for CIFAR-10; 2 for ImageNet.
+  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
+Table 9: Model architectures.
+      Dataset       Architecture
+      CIFAR-10       WideResNet-94-16 [6]
+      RaWideResNet-70-16 [35]
+      WideResNet-34-20 [32]
+      XCiT-L12 [13]
+      PreActResNet-18 [19] (backbone)
+      ResNeXt29-32x4d [21] (classifier)
+      ImageNet       Swin-L [29]
+      ConvNeXt-L+ConvStem [44] (backbone)
+      DeiT Base [45] (classifier)
+Appendix F Related Works
 Existing adversarial defenses are divided into three categories: training-time, test-time, and preemptive.
@@
 It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
-Appendix B Algorithms of Fast Preemption and
-Preemptive Reversion
-The complete defense algorithm of Fast Preemption is presented in Algorithm 1.
-With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
-The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
-Algorithm 1 Fast Preemption
-1:Original image $x$, classifier $f_{c}$, backbone model $f_{b}$, maximum $L_{\infty}$-norm distance $\epsilon$, learning rate $\eta$, maximum forward steps $T_{\rightarrow}$ and backward steps $T_{\leftarrow}$, number of examples for averaging perturbations $N$.
-2:Robust example $x^{r}$.
-3:// Forward Propagation
-4:$x_{0}=x$
-5:for $t<T_{\rightarrow}$ do
-6:       for $n=1,...,N$ do
-7:             $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
-8:             $x_{t,n}=x_{t}+\delta_{n}$
-9:             $x_{t,n}^{r}=\prod_{x,\epsilon}(x_{t,n}-\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM decreases $\ell$
-10:       end for
-11:       $\delta_{t}^{r}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{r}-x_{t}$$\triangleright$ averages perturbations
-12:       $x_{t+1}=x_{t}+\eta\cdot\delta_{t}^{r}$$\triangleright$ direct enhancement
-13:end for
-14:// Backward Propagation
-15:$x_{0}=x_{T_{\rightarrow}}$
-16:for $t<T_{\leftarrow}$ do
-17:       for $n=1,...,N$ do
-18:             $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
-19:             $x_{t,n}=x_{t}+\delta_{n}$
-20:             $x_{t,n}^{a}=\prod_{x,\epsilon}(x_{t,n}+\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM increases $\ell$
-21:       end for
-22:       $\delta_{t}^{a}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{a}-x_{t}$$\triangleright$ averages perturbations
-23:       $x_{t+1}=x_{t}-\eta\cdot\delta_{t}^{a}$$\triangleright$ reverses the attack
-24:end for
-25:return $x^{r}=x_{T_{\leftarrow}}$
-Algorithm 2 Preemptive Reversion
-1:Robust example $x^{r}$.
-2:Reversed example $\tilde{x}$.
-3:$(x^{r})^{r}=PreemptiveDefense(x^{r})$$\triangleright$ secondary defense
-4:$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r})$$\triangleright$ reverses the secondary defense
-5:return $\tilde{x}$
-Table 3: Experiment Settings.
-       Type        Approach        Settings
-               Adversarial        Attack        PGD [31]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
-       AutoAttack [11]
-       CW [7]
-       DeCoW [28]
-       MCG [53]        Maximum Queries = 20,000
-               Reversion        Attack        JPEG        Quality Factor = 50%
-       DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
-       Fast Preemption (ours)        Same with our defense
-       Defense        DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
-       Bi-Level [32]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
-       $A^{5}$ [17]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
-       Fast Preemption (ours)        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
-  • 1 for CIFAR-10; 2 for ImageNet.
-  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
-Appendix C Experiment Settings
-The full experimental settings used for evaluations are listed in Tab. 3.
-Appendix D Model Architectures
-We evaluated attacks and defenses using models with various architectures, as listed in Tab. 4.
-Table 4: Model architectures.
-      Dataset       Architecture
-      CIFAR-10       WideResNet-94-16 [6]
-      RaWideResNet-70-16 [35]
-      WideResNet-34-20 [32]
-      XCiT-L12 [13]
-      PreActResNet-18 [19] (backbone)
-      ResNeXt29-32x4d [21] (classifier)
-      ImageNet       Swin-L [29]
-      ConvNeXt-L+ConvStem [44] (backbone)
-      DeiT Base [45] (classifier)
-Appendix E Implementation
-Our implementation partially leveraged the ART [33] and RobustBench [12] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
-Appendix F Ablation Studies
-We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
-The attacks were PGD [31] and AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-As illustrated in Fig.
-6 of Sec.
-7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [32] in terms of clean and robust accuracies and time cost.
-These studies and results were aimed at identifying the optimal strategies.
-Table 5: Performance with different classifiers.
-       Ablation        Time Cost(s)        Transferable Defense
-       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
-       Ground Truths        100        28        97.2        86.9        83.1
-       ResNeXt        95.9        28        95.4        85.7        82.7
-       Bi-Level        89.1        29        91.9        83.9        81.0
-  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [32].
-The transferable defenses were all evaluated using RaWideResNet-70-16 [35] as the victim model.
-Bold indicates the best performance.
-F.1 Classifier
-In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction.
-As demonstrated in Tab.
-5, a classifier with higher accuracy leads to better performance.
-As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning.
-Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction.
-Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
-Table 6: Performance of different backbone models.
-Backbone Model: Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
-Backbone Robustness Clean (%) 89.1 93.9 88.1
-PGD (%) 43.9 77.1 61.8
-AutoAttack (%) 42.4 74.0 58.6
-Time Cost (s) for Defense 28 69 26
-White-Box Defense Clean (%) 95.5 95.8 93.8
-PGD (%) 84.9 91.6 85.9
-AutoAttack (%) 81.8 90.2 83.3
-Transferable Defense Victim: RaWideResNet-70-16 [35] Clean (%) 95.4 95.5 95.3
-PGD (%) 85.7 89.7 87.6
-AutoAttack (%) 82.7 88.0 85.9
-Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
-PGD (%) 72.7 78.6 79.6
-AutoAttack (%) 70.8 75.7 77.3
-  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [21] as the classifier model, and 100 steps backward propagation as Bi-Level [32].
-The transferable defenses were all evaluated using RaWideResNet-70-16 [35] and XCiT-L12 [13] as the victim models, different from the backbone models.
-Bold indicates the best performance.
-F.2 Backbone Model
-The backbone model determines the defense time cost and performance.
-As illustrated in Fig.
-7, different backbone models learn different features.
-To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec.
-1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [39].
-Figure 7: Different backbone models learn different features.
-PreActResNet-18 [19], WideResNet-94-16 [6], and Bi-Level [32] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
-As shown in Tab.
-6, the best adversarially trained model was WideResNet-94-16 [6], so it is not surprising that its white-box defense outperformed that of the other backbone models.
-However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [13]), and the time cost was much higher due to its complexity.
-We thus empirically selected PreActResNet-18 [19] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [6].
-Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
-F.3 Forward-Backward Cascade Learning
-As observed in Fig.
-8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0).
-In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90).
-Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
-Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig.
-9.
-The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs.
-F0-B3).
-Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
-Figure 8: Forward propagation (F) converges faster while backward propagation (B) alleviates overfitting.
-The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
-The numbers in brackets represent robust accuracy against AutoAttack (Standard).
-“Alternate” represents the running of forward and backward propagation alternatively for 100 steps in total.
-Figure 9: By combining forward (F) and backward (B) propagation, Fast Preemption obtains performance comparable to that of the SOTA preemptive defense [32] by using only three steps (F1–B2).
-The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
-The numbers in brackets represent robust accuracy against AutoAttack (Standard).
-Table 7: Performance for different $N$ and optimization algorithms.
-       Ablation        Time Cost (s)        Transferable Defense
-       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
-       PGD        1        0.7        95.3        87.5        85.8
-       FGSM        1        0.04        95.4        86.8        85.4
-       PGD        20        0.8        95.2        87.7        85.9
-       FGSM        20        0.04        95.5        87.8        85.9
-  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [35] as the victim model. Bold indicates the best performance.
-F.4 Optimization Algorithm
-The results in Tab.
-7 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost.
-Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
-Appendix G Evaluation across More Models
-In the main paper, we concentrated on comparisons with top-performing adversarial training methods [6, 29].
-In this section, we extend our evaluation to include defenses tested across nine additional victim models on CIFAR-10 [26] from RobustBench [12] in a transferable context.
-These models encompass diverse architectures representative of platforms such as Twitter and Facebook, differing from both the classifier [21] and the backbone model [19].
-As shown in Tab.
-8, our method demonstrates significant performance gains compared to no proactive defense and the SOTA proactive defense [32].
-Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively.
-Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
-Table 8: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-All proactive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
-Model Metric No Proactive Defense Bi-Level [32] Fast Preemption (ours)
-Defense Learning Backbone N/A Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
-Time Cost (s) N/A 28 0.35 0.04
-RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
-Robust 71.4 81.3 86.8 85.9
-WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
-Robust 71.2 80.9 86.4 86.0
-WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
-Robust 62.1 76.2 78.7 80.8
-WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
-Robust 64.8 77.7 81.3 83.2
-PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
-Robust 42.4 54.7 55.4 61.2
-ResNest152 [41] Clean 86.9 87.6 91.2 92.4
-Robust 62.4 73.2 75.8 78.2
-ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
-Robust 52.0 68.1 66.7 72.0
-ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
-Robust 51.4 67.5 67.8 72.8
-XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
-Robust 57.9 71.3 74.6 76.0
-Average Clean 89.9 89.8 93.4 94.1
-Robust 59.5 72.3 74.8 77.3
-  • Bold indicates the best performance.
-Appendix H Defense against More Attacks
-Our primary evaluation adheres to the RobustBench [12] protocol, employing the strongest attack, AutoAttack [11].
-As detailed in Sec.
-5.2, this includes tests with standard and randomized AutoAttack variants, encompassing white-box APGD [11], FAB [10], and black-box Square [4] attacks, alongside EOT settings [5].
-In this section, we extend the evaluation to include three additional attacks: white-box Carlini-Wagner (CW) [7], black-box Deformation-Constrained Warping (DeCoW) [28], and black-box Meta Conditional Generator (MCG) [53], demonstrating the broad applicability and generalizability of our defense against diverse attack strategies.
-As shown in Tab.
-9, our defense achieves robust accuracy of 90.3%, 93.6%, and 87.3% against CW, DeCoW, and MCG attacks, respectively, outperforming SOTA adversarial training [6] and proactive defenses [32].
-Notably, these accuracies exceed those obtained against AutoAttack, further highlighting the robustness and reliability of our primary evaluation.
-Table 9: Robust accuracy (%) against various adversarial attacks with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-      Threat Model       Attack       WideResNet-94-16 [6]       Bi-Level [32]       Fast Preemption (ours)
-      White-Box       AutoAttack (Standard) [11]       74.0       82.1       86.3
-      CW [7]       80.8       85.4       90.3
-      Black-Box       DeCoW [28]       88.8       89.7       93.6
-      MCG [53]       76.1       83.3       87.3
-  • Bold indicates the best performance.
-Appendix I Visual Quality
-The noise budgets utilized (as detailed in Tab.
-3) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection.
-As illustrated in Fig.
-10, for 1,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images.
-Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
-Figure 10: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA proactive defense [32].
-All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images.
-SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
-I.1 Structural Similarity Index (SSIM)
-The SSIM is a metric for measuring similarity between two images.
-It considers luminance, contrast, and structural information.
-The SSIM between two images $x$ (reference) and $y$ (distorted) is defined as:
-$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (15)
-where:
-  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
-  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
-  • $\sigma_{xy}$: Covariance between $x$ and $y$.
-  • $C_{1}$ and $C_{2}$: Small constants added to stabilize the metric in the presence of weak denominators, defined as:
-    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
-    where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit images), and $K_{1}$ and $K_{2}$ are small positive constants (commonly, $K_{1}=0.01$ and $K_{2}=0.03$).
-The SSIM value ranges from -1 to 1, where:
-  • 1: Perfect similarity.
-  • 0: No similarity.
-  • Negative values indicate dissimilarity.
-SSIM is typically calculated locally using a sliding window across the image, and the overall SSIM is the average of local SSIM values.
-Higher SSIM values denote subtler deviations from the reference images.
-I.2 Learned Perceptual Image Patch Similarity (LPIPS)
-The LPIPS is a perceptual similarity measure that evaluates the perceptual distance between two images.
-It uses deep neural network features to compare the visual content and correlates better with human perception than traditional metrics like SSIM.
-The LPIPS distance between two images $x$ (reference) and $y$ (distorted) is computed as:
-$LPIPS(x,y)=$ (17)
-$\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}w_{l}\|\hat{\phi}_{l}(x)_{h,w}-\hat{\phi}_{l}(y)_{h,w}\|_{2}^{2},$
-where:
-  • $l$: Layer index in the deep network.
-  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
-  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
-  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
-  • $w_{l}$: Learned scalar weights for each layer $l$ (trained to match human judgments).
-LPIPS uses a pretrained deep neural network to extract feature maps from intermediate layers.
-The metric normalizes the feature maps and computes the difference between them, weighted according to the learned weights $w_{l}$.
-The LPIPS value typically ranges from 00 (perfect perceptual similarity) to higher values indicating greater dissimilarity.
-LPIPS provides a perceptually grounded comparison, making it especially useful for evaluating subtle changes in visual quality.

### v7

@@
-Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense
+Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Preemptive Adversarial Defense
 Abstract
-Deep learning has advanced significantly but remains vulnerable to adversarial attacks, compromising its reliability.
-While conventional defenses typically mitigate perturbations post-attack, few studies explore proactive strategies for preemptive protection.
-This paper proposes Fast Preemption, a novel defense that neutralizes adversarial effects before third-party attacks.
-By employing distinct models for input labeling and feature extraction, Fast Preemption enables an efficient, transferable preemptive defense with state-of-the-art robustness across diverse systems.
-To further enhance efficiency, we introduce a forward-backward cascade learning algorithm that generates protective perturbations, leveraging forward propagation for rapid convergence and iterative backward propagation to mitigate overfitting.
-Executing in just three iterations, Fast Preemption surpasses existing training-time, test-time, and preemptive defenses.
-Additionally, we propose the first effective white-box adaptive reversion attack to evaluate the reversibility of preemptive defenses, demonstrating that our approach remains secure unless the backbone model, algorithm, and settings are entirely compromised.
-This work establishes a new paradigm for proactive adversarial defense.
+Deep learning has made significant strides but remains susceptible to adversarial attacks, undermining its reliability.
+Most existing research addresses these threats after attacks happen.
+A growing direction explores preemptive defenses like mitigating adversarial threats proactively, offering improved robustness but at cost of efficiency and transferability.
+This paper introduces Fast Preemption, a novel preemptive adversarial defense that overcomes efficiency challenges while achieving state-of-the-art robustness and transferability, requiring no prior knowledge of attacks and target models.
+We propose a forward-backward cascade learning algorithm, which generates protective perturbations by combining forward propagation for rapid convergence with iterative backward propagation to prevent overfitting.
+Executing in just three iterations, Fast Preemption outperforms existing training-time, test-time, and preemptive defenses.
+Additionally, we introduce an adaptive reversion attack to assess the reliability of preemptive defenses, demonstrating that our approach remains secure in realistic attack scenarios.
 1 Introduction
@@
 Defending against such attacks remains a critical yet unresolved challenge.
-Passive defenses [6, 29, 34, 22] enhance robustness at the cost of classification accuracy [54] and remain susceptible to adaptive attacks [7, 46].
-Proactive defenses [32, 17], while mitigating threats preemptively, often lack real-time efficiency and require paired protection-inference modules, limiting transferability.
-Moreover, users relying on third-party models must contend with potential backdoor threats [39].
-Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
-To address these challenges, we propose a proactive defense strategy, Fast Preemption, as illustrated in Fig.
+Training-time and test-time defenses [6, 26, 31, 22] mitigate adversarial threats after attacks happen, improving robustness but reducing classification accuracy [51] and remaining vulnerable to adaptive attacks [7, 43].
+In contrast, preemptive defenses [29, 17] preemptively modify media to neutralize future adversarial manipulations.
+However, preemptive defenses suffer from inefficiency and rely on paired protection-inference modules, limiting transferability.
+Additionally, the protection modules in preemptive defenses are dependent on the paired, provided models, rendering them susceptible to potential backdoor threats [36].
+Figure 1: The proposed Fast Preemption offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
+To address these challenges, we propose Fast Preemption, a novel preemptive defense illustrated in Fig.
 1.
-Fast Preemption preemptively “attacks” media on the user side to enhance robustness before potential third-party attacks.
-Specifically, it learns robust features, predicts possible adversarial perturbations, and then amplifies these features while reversing the predicted perturbations for protection.
-As a result, any subsequent adversarial attack is inherently neutralized.
-This approach offers three key advantages.
-(i) Unlike passive defenses, Fast Preemption enhances threat resistance without compromising classification accuracy and requires no modifications to the classification system.
-(ii) It supports plug-and-play deployment without paired protection-inference modules, mitigating security and privacy concerns associated with external model providers [16, 39].
-(iii) Traditional training- and test-time defenses, integrated within classification models, risk full-system compromise if breached.
-In contrast, Fast Preemption localizes vulnerabilities to individual users, as each employs distinct configurations.
-Thus, Fast Preemption satisfies four fundamental requirements:
-  • Preserve clean accuracy and improve robust accuracy: Basic requirements for all defenses.
-  • Not require ground truths: It is impractical to label each input manually.
-  • Provide an efficient defense: Time cost per process must be satisfactory for user interaction.
-  • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
-We introduce Fast Preemption, a preemptive defense comprising three key components: a classifier model for automatic input labeling, a backbone model (distinct from the classifier) for feature learning, and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations.
-We evaluate Fast Preemption on two datasets and thirteen models, defending against five representative adversarial attacks.
-Results demonstrate its superiority in clean and robust accuracy, transferability, and computational efficiency over benchmark training-time, test-time, and preemptive defenses [6, 29, 34, 32, 17].
-For example, on CIFAR-10 [26], Fast Preemption outperformed the state-of-the-art (SOTA) preemptive defense [32], achieving 2.5% and 5.1% higher clean and robust accuracy, respectively, against AutoAttack (Standard) [11], while being 700 times faster.
-Shannon famously stated, “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them” [43].
-Following this principle, Bi-Level [32] first introduced a reversion attack but failed, asserting their defense was irreversible.
-In contrast, our reversion attack successfully circumvented their defense, making ours the first effective adaptive reversion attack in the white-box setting, termed Preemptive Reversion, to assess the reversibility of preemptive defenses.
-Preemptive Reversion applies a secondary preemptive defense to the already protected example and then reverses the newly introduced perturbations.
-Since the secondary defense generates perturbations similar to the first under identical conditions, this process effectively tests a defense’s resilience to reversion.
-Additionally, we propose a more rigorous evaluation protocol, which mandates a decrease in clean accuracy after preemptive defense (achieved by selecting a suboptimal classifier).
-If clean accuracy improves post-reversion, the attack is successful; if it further deteriorates due to distortion, reversion has failed.
-Results of reversion attacks show that Preemptive Reversion effectively removes protective perturbations in a white-box setting, where all defense details are exposed.
-In contrast, blind JPEG distortion and adversarial purification [34] further degrade the image rather than neutralizing protections.
-However, this white-box assumption is unrealistic for real-world applications of Fast Preemption, as the framework allows users to customize their defense settings individually.
-Beyond the white-box scenario, existing reversion methods fail to circumvent Fast Preemption’s protection.
-We summarize our contributions as follows: (i) We introduce Fast Preemption, a novel preemptive defense that enhances transferability without requiring ground truths or paired modules.
-It is the first approach to utilize distinct models for the classifier and backbone.
-(ii) We propose the forward-backward (Fore-Back) cascade learning algorithm, which accelerates preemptive defense, enabling efficient user interaction.
-(iii) We develop Preemptive Reversion, the first effective adaptive reversion attack for assessing the robustness of preemptive defenses in a white-box setting.
-Additionally, we establish a more comprehensive evaluation protocol to determine reversion success.
+Fast Preemption learns robust features, predicts potential adversarial perturbations, and amplifies these features while reversing the predicted perturbations, effectively neutralizing subsequent attacks.
+It offers six key advantages over other defenses:
+  • No reliance on ground truths: While ground-truth labels offer the highest protection, manual labeling is less practical. Fast Preemption operates without this requirement.
+  • Plug-and-play: Unlike training-time defenses, Fast Preemption requires no modifications to the classification system.
+Unlike existing preemptive defenses, it eliminates reliance on paired protection-inference modules, reducing security and privacy risks associated with external model providers [16, 36].
+  • Transferable: Fast Preemption provides generalized protection without requiring prior knowledge of target models or attack algorithms.
+  • Efficient: A novel learning algorithm greatly improves efficiency compared to existing preemptive defenses.
+  • State-of-the-art (SOTA) clean and robust accuracy: Unlike training-time and test-time defenses, Fast Preemption enhances threat resistance without compromising classification accuracy.
+It achieves the best clean and robust accuracy among existing adversarial defenses.
+  • Resilient to adaptive attacks: Fast Preemption allows users to set protection strength and select a backbone model independently of the target model, enhancing resilience to adaptive attacks beyond existing test-time and preemptive defenses.
+To achieve these objectives, we design Fast Preemption with three key components: a user-configurable classifier for automatic input labeling, a user-configurable backbone model (which may differ from the classifier) for feature learning, and a forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations.
+The Fore-Back cascade learning algorithm integrates forward and backward propagation to overcome their individual limitations.
+Forward propagation ensures rapid convergence but overfits to the backbone model, reducing transferability.
+Conversely, backward propagation mitigates overfitting but requires many iterations, making it computationally expensive.
+By combining both, Fore-Back achieves efficient and robust protection.
+We evaluate Fast Preemption on two datasets and thirteen models, defending against four representative adversarial attacks.
+Results demonstrate its superiority in clean and robust accuracy, transferability, and computational efficiency over benchmark training-time, test-time, and preemptive defenses.
+For example, on CIFAR-10 [23], Fast Preemption outperforms the SOTA preemptive defense [29], achieving 2.5% and 5.1% higher clean and robust accuracy, respectively, against AutoAttack (Standard) [11], while being 700 times faster.
+We further explore stronger adaptive attack scenarios where the attacker has full knowledge of our defense and attempts to reverse its protective perturbations.
+Moon et al.
+[29] introduced the first reversion attack but failed, claiming their preemptive defense was irreversible.
+We propose Preemptive Reversion, a more powerful reversion attack that successfully bypasses their defense in a white-box setting (i.e., both the defense algorithms and settings have been compromised).
+Preemptive Reversion applies a secondary preemptive defense to an already protected example and then reverses the newly introduced perturbations.
+Since the secondary defense generates perturbations similar to the first under identical conditions, this process effectively evaluates a defense’s susceptibility to reversion.
+We employ Preemptive Reversion alongside methods such as blind JPEG distortion and adversarial purification [31] to assess the reliability of Fast Preemption against adaptive attacks.
+Since Fast Preemption allows for user-configurable parameterization (e.g., selection of the backbone model), and these configurations remain confidential, all adaptive attacks fail to bypass its protection, highlighting its robustness.
+However, in the less-realistic scenario where the user configurations are compromised, only Preemptive Reversion successfully removes the protective perturbations without introducing significant distortion, demonstrating the aggressiveness of the proposed attack.
+To our knowledge, Preemptive Reversion remains the only viable approach for evaluating the reliability of preemptive defenses against adaptive attacks.
+We summarize our contributions as follows: (i) We introduce Fast Preemption, a more effective plug-and-play preemptive defense that requires neither ground truths nor prior knowledge of attacks or target models, while also providing resistance to adaptive attacks.
+(ii) We propose the Fore-Back cascade learning algorithm, which significantly enhances the transferability and learning speed of preemptive defenses.
+(iii) We develop Preemptive Reversion, a more robust tool for evaluating the reliability of preemptive defenses against adaptive attacks.
+Figure 2: Fast Preemption framework.
+The classifier labels inputs automatically, which improves clean accuracy and thereby improves defense performance.
+The backbone model learns crucial features, so it must be adversarially trained.
+The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
 2 Related Works
-Existing adversarial defenses can be classified into three main categories: training-time, test-time, and preemptive defenses. A detailed review is provided in Appendix F.
+Existing adversarial defenses can be classified into three main categories: training-time, test-time, and preemptive defenses.
 Training-time defenses aim to develop robust deep learning models that can correctly classify adversarial examples (AEs) based on their ground truth labels [18, 39, 38, 19, 13, 47, 14, 32, 51, 6, 26, 40].
@@
 Second, training-time defenses often sacrifice clean accuracy to improve robust accuracy.
-Test-time defenses are typically divided into detection and purification.
+Test-time defenses are divided into confusion, detection, and purification.
+Confusion methods integrate multiple layers or models, so that if an attacker bypasses one, other layers can resist the threat [42, 3].
+However, Alfarra et al.
+[42] and Wang et al.
+[44] acknowledge that this strategy is vulnerable when the attacker is aware of the defense and targets all layers in a single attack.
 Detection models identify AEs by distinguishing them from benign inputs [9, 53, 24, 22], but detection is challenging for unseen attacks, as detectors are often trained on known AEs.
@@
 A key limitation of purification is that the difference between original and generated examples can degrade clean accuracy, and counterattack potential remains [7, 43].
-Figure 2: Fast Preemption framework.
-The classifier labels inputs automatically, which improves clean accuracy and thereby improves defense performance.
-The backbone model learns crucial features, so it must be adversarially trained.
-The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
 Preemptive defenses aim to generate robust examples (REs) that replace the original images, which are then discarded [37, 29, 17].
@@
 However, these approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, reduced robustness compared to other defense mechanisms.
-Alfarra et al.
-[3] proposed a test-time defense incorporating preemptive protection as an anti-adversary layer guided by a standard classifier.
-However, their method is inherently vulnerable if the classifier is compromised, a limitation they explicitly acknowledge.
-This weakness undermines its practical viability in test-time scenarios.
-While they describe this setting as the “least realistic,” we argue that it aligns with the standard white-box attack model, where adversaries have full access to the classifier.
-Under even stricter assumptions, adversaries may also be aware of the defense mechanism itself (white-box reversion).
-Moreover, their approach relies solely on forward propagation learning, increasing susceptibility to overfitting.
 3 Preemptive Defense: Fast Preemption
@@
 Classifier: Correct input labels determines the correct optimization direction for defense.
-However, for efficient user interaction, it is impractical to label each input manually.
+However, for efficient user interaction, it is less practical to label each input manually.
 Classifier $f_{c}$ labels these inputs automatically.
@@
 In summary, the forward and backward propagation both enhance robustness, but in different optimization directions.
-The complete algorithm is presented in Algorithm 1 of Appendix D.
+The complete algorithm is presented in Algorithm 1 of Appendix A.
 3.3.1 Loss Function
@@
 $s.t.\ ||x_{*}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
-where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
+where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Eq. 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
 Whereas backward propagation requires many steps to converge, forward propagation converges quickly.
@@
 4 Adaptive Attack: Preemptive Reversion
-A reversion attack is used to assess the reversibility of a preemptive defense.
-While Bi-Level [32] first proposed a reversion attack but failed to reverse their defense, claiming it was irreversible, our reversion attack successfully overcame their defense, making ours the first effective reversion attack.
+A reversion attack is used to assess the reliability of a preemptive defense.
+While Moon et al.
+[29] first proposed a reversion attack against their proposed defense, Bi-Level, but failed to reverse it, claiming it was irreversible, we propose a stronger reversion attack successfully overcame Bi-Level.
 In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks.
-Preemptive Reversion is the only viable approach to reverse preemptive defenses in the white-box scenario.
-However, without full knowledge of the backbone model and settings, our Fast Preemption defense remains difficult to reverse, ensuring its reliability.
 Figure 3: Illustration of Preemptive Reversion algorithm.
@@
 5).
-On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig.
+On the basis of this finding, we devised the adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig.
 3.
-The complete algorithm is presented in Algorithm 2 of Appendix D.
+The complete algorithm is presented in Algorithm 2 of Appendix A.
 Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$.
@@
   • White-Box: The attacker has complete knowledge of both the backbone model and defense algorithm (including settings), enabling them to replicate the secondary defense.
-This scenario, while the strongest, is impractical due to varying user settings.
+This scenario, while the strongest, is less realistic due to varying user settings.
   • Black-Box: The attacker knows the defense algorithm (typically public) but lacks knowledge of the backbone model, which can vary across users.
 The challenge in this scenario is that the features learned in the secondary defense may differ from those learned by the user.
-Our proposed Preemptive Reversion attack is the first effective method for evaluating the reversibility of preemptive defenses, including Fast Preemption, but it is only feasible in the white-box scenario.
+Our proposed Preemptive Reversion attack is effective for evaluating the reversibility of preemptive defenses, including Fast Preemption, but it is only feasible in the white-box scenario.
 Without full knowledge of the defense, Fast Preemption remains reliable.
@@
 The model architectures are listed in Tab.
-9 of Appendix E.
+5 of Appendix B.
 Adversarial attacks.
-We evaluated defenses against five adversarial attacks: white-box PGD [31], AutoAttack [11], and Carlini-Wagner (CW) attack [7]; black-box Deformation-Constrained Warping (DeCoW) [28] and Meta Conditional Generator (MCG) [53].
+We evaluated defenses against four adversarial attacks: white-box attacks include AutoAttack [11] and Carlini-Wagner (CW) [7]; black-box attacks include Deformation-Constrained Warping (DeCoW) [25] and Meta Conditional Generator (MCG) [50].
 For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, FAB [10], and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]).
@@
 6 of Appendix C.
-Test-time defenses. We used DiffPure [34] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [29] depending on the dataset.
-Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [32], and the latest preemptive defense, $A^{5}$ [17].
-5.1 Evaluation metrics
+Test-time defenses.
+We used DiffPure [31] as the benchmark test-time defense.
+The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [26] depending on the dataset.
+Preemptive defenses.
+We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [29], and the latest preemptive defense, $A^{5}$ [17].
+Evaluation metrics.
 We evaluated performance using clean accuracy and robust accuracy, where higher values indicate better defense effectiveness.
@@
 8
-Fast Preemption (ours) + DiffPure [34] 0.34 94.2 89.1 92.8 89.7
+Combination Ours+[31] 0.34 94.2 89.1 92.8 89.7
 ImageNet Training-Time Swin-L [26] 0 93.0 67.0 59.5 67.5
@@
 8
-Fast Preemption (ours) + DiffPure [34] 2.5 93.3 82.9 87.6 83.4
+Combination Ours+[31] 2.5 93.3 82.9 87.6 83.4
   • Bold indicates the best performance.
@@
 6 Performance
-The experiments were conducted using the settings outlined in Tab. 8 of Appendix E. Time costs were recorded on an NVIDIA A100 40GB GPU.
+The experiments were conducted using the settings outlined in Tab. 4 of Appendix B. Time costs were recorded on an NVIDIA A100 40GB GPU.
 Table 3: Preemptive defenses resist reversion attacks.
@@
   • Numbers in red demonstrate that only our reversion attack is feasible but only in the white-box setting.
-Figure 4: Perturbations of adversarial attacks and the proposed preemptive defenses.
-Visual comparison between the original image and the RE demonstrates that Fast Preemption perturbations are imperceptible.
-Additionally, comparison of perturbations between AutoAttack and Fast Preemption reveals that preemptive defense is more likely to succeed when perturbations are similar.
 6.1 Robustness against Adversarial Attacks
-We evaluated transferable defense performance under the condition where the defense lacks knowledge of both the attacks and victim models [6, 29].
-Table 1 and Tab.
+We evaluated transferable defense performance under the condition, where the defense lacks knowledge of both the attacks and victim models [6, 26].
+Tab.
+1 and Tab.
 2 compares Fast Preemption with benchmark defenses against four attacks.
@@
 1.
-Figure 4 provides a qualitative analysis, comparing the original images with their corresponding REs.
+Figure 4: Perturbations of adversarial attacks and preemptive defenses.
+Visual comparison between the original image and the robust example (RE) shows that Fast Preemption perturbations are imperceptible.
+Furthermore, comparison of perturbations between AutoAttack and Fast Preemption indicates that the preemptive defense is more likely to succeed when perturbations are similar.
+Fig.
+4 provides a qualitative analysis, comparing the original images with their corresponding REs.
 The perturbations introduced by Fast Preemption are barely perceptible.
-The average SSIM and LPIPS scores for 2,000 examples from CIFAR-10 and ImageNet are 0.97 and 0.01, respectively, both surpassing those of standard AutoAttack and the SOTA preemptive defense Bi-Level [32] (more details in Appendix A).
+The average SSIM and LPIPS scores for 2,000 examples from CIFAR-10 and ImageNet are 0.97 and 0.01, respectively, both surpassing those of AutoAttack and the SOTA preemptive defense Bi-Level [29] (more details in Appendix D).
 Additionally, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations closely resemble those of the attack.
@@
 Other reversion attacks failed, further degrading clean accuracy due to distortion.
-Notably, Preemptive Reversion is only effective in the white-box setting, as Fast Preemption remains irreversible without full knowledge of the defense.
-Figure 5 corroborates our findings, showing that the white-box reversion successfully eliminates most protective perturbations, whereas the black-box reversion fails.
+Notably, Preemptive Reversion is effective only in the white-box setting, which is less realistic for our preemptive defense, as our defense allows confidential user configurations (e.g., selection of the backbone model).
+In the black-box setting, where the backbone models are unknown, all adaptive attacks fail to bypass the preemptive defense, further demonstrating its reliability.
+In contrast, other preemptive defenses, such as Bi-Level [29], rely on paired protection-inference models, meaning that compromising the target model is sufficient to reverse user protections.
+Fig.
+5 corroborates our findings, showing that the white-box reversion successfully eliminates most protective perturbations, whereas the black-box reversion fails.
 However, due to the randomization inherent in Fast Preemption (i.e., randomly initialized FGSM), the white-box reversion cannot entirely remove all protective perturbations.
@@
 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [29], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to 0.04 seconds.
-These studies were conducted to identify optimal configurations, with further ablation results provided in Appendix B.
+These studies were conducted to identify optimal configurations, with further ablation results provided in Appendix E.
 Figure 6: Compared with the SOTA preemptive defense Bi-Level [29], our framework improved the clean and robust accuracies, and our learning algorithm reduced the time cost for defense.
 8 Conclusion
-We propose Fast Preemption, an efficient and transferable preemptive adversarial defense that introduces a new layer in the defense hierarchy: Proactive Defense (ours) → Training-Time Defense → Test-Time Defense.
-Unlike conventional training- and test-time defenses, Fast Preemption is plug-and-play, transferable, and requires no prior knowledge, enabling users to safeguard media independently without relying on third-party interventions.
-Our approach also addresses the efficiency limitations of existing proactive defenses, achieving processing speeds 700 times faster than prior methods, an essential improvement for real-world applications, where significant delays (e.g., 28 seconds per image [32]) are impractical.
-Moreover, Fast Preemption maintains or improves clean accuracy while achieving SOTA robust accuracy, overcoming the common trade-off where defenses degrade clean accuracy.
-Despite its advantages, our defense exhibits limitations against the proposed reversion attack in the white-box setting.
-While allowing users to customize the backbone model as a user-side key mitigates reversion attacks, further work is needed to simplify user requirements.
-Future research should also investigate the relationship between backbone models and transferability, as discussed in Sec.
-B.2.
+We propose Fast Preemption, an efficient, transferable, plug-and-play preemptive adversarial defense that introduces a new layer in the defense hierarchy: Preemptive Defense (ours) → Training-Time Defense → Test-Time Defense.
+Fast Preemption minimizes the need for prior knowledge, allowing users to independently safeguard media without relying on third-party interventions.
+It achieves SOTA clean and robust accuracy among training-time, test-time, and preemptive defenses while addressing the efficiency limitations of existing preemptive approaches, achieving processing speeds up to 700 times faster.
+Additionally, we propose Preemptive Reversion, a more robust reversion attack that demonstrates our preemptive defense’s resilience against reversion by incorporating a strategy that maintains the confidentiality of user configurations.
+However, our defense exhibits limitations against the proposed reversion attack.
+While allowing users to customize defense configurations mitigates reversion attacks, further work aims to simplify user requirements.
+Future research should also investigate the relationship between backbone models and transferability.
 Supplementary Material
-This supplementary material provides a comprehensive analysis of visual quality (Appendix A) and presents an extensive ablation study (Appendix B), investigating critical components such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
-Additionally, it includes evaluation results against a broader set of models (Appendix C), elaborates on the algorithms underlying the proposed adversarial defense and adaptive reversion attack (Appendix D), and details the experimental configurations for both attack and defense scenarios (Appendix E).
-Appendix A Visual Quality
-The noise budgets utilized (as detailed in Tab.
-8) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection.
-As illustrated in Fig.
-7, for 2,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images.
-Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
-Figure 7: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA proactive defense [32].
-All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images.
-SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
-A.1 Structural Similarity Index (SSIM)
-The SSIM is a metric for measuring similarity between two images.
-It considers luminance, contrast, and structural information.
-The SSIM between two images $x$ (reference) and $y$ (distorted) is defined as:
-$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (15)
-where:
-  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
-  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
-  • $\sigma_{xy}$: Covariance between $x$ and $y$.
-  • $C_{1}$ and $C_{2}$: Small constants added to stabilize the metric in the presence of weak denominators, defined as:
-    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
-    where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit images), and $K_{1}$ and $K_{2}$ are small positive constants (commonly, $K_{1}=0.01$ and $K_{2}=0.03$).
-The SSIM value ranges from -1 to 1, where:
-  • 1: Perfect similarity.
-  • 0: No similarity.
-  • Negative values indicate dissimilarity.
-SSIM is typically calculated locally using a sliding window across the image, and the overall SSIM is the average of local SSIM values.
-Higher SSIM values denote subtler deviations from the reference images.
-A.2 Learned Perceptual Image Patch Similarity (LPIPS)
-The LPIPS is a perceptual similarity measure that evaluates the perceptual distance between two images.
-It uses deep neural network features to compare the visual content and correlates better with human perception than traditional metrics like SSIM.
-The LPIPS distance between two images $x$ (reference) and $y$ (distorted) is computed as:
-$LPIPS(x,y)=$ (17)
-$\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}w_{l}\|\hat{\phi}_{l}(x)_{h,w}-\hat{\phi}_{l}(y)_{h,w}\|_{2}^{2},$
-where:
-  • $l$: Layer index in the deep network.
-  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
-  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
-  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
-  • $w_{l}$: Learned scalar weights for each layer $l$ (trained to match human judgments).
-LPIPS uses a pretrained deep neural network to extract feature maps from intermediate layers.
-The metric normalizes the feature maps and computes the difference between them, weighted according to the learned weights $w_{l}$.
-The LPIPS value typically ranges from 00 (perfect perceptual similarity) to higher values indicating greater dissimilarity.
-LPIPS provides a perceptually grounded comparison, making it especially useful for evaluating subtle changes in visual quality.
-Appendix B Ablation Studies
-We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
-The attacks were PGD [31] and AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-As illustrated in Fig.
-6 of Sec.
-7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [32] in terms of clean and robust accuracies and time cost.
-These studies and results were aimed at identifying the optimal strategies.
-Table 4: Performance with different classifiers.
-       Ablation        Time Cost(s)        Transferable Defense
-       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
-       Ground Truths        100        28        97.2        86.9        83.1
-       ResNeXt        95.9        28        95.4        85.7        82.7
-       Bi-Level        89.1        29        91.9        83.9        81.0
-  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [32].
-The transferable defenses were all evaluated using RaWideResNet-70-16 [35] as the victim model.
-Bold indicates the best performance.
-B.1 Classifier
-In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction.
-As demonstrated in Tab.
-4, a classifier with higher accuracy leads to better performance.
-As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning.
-Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction.
-Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
-Table 5: Performance of different backbone models.
-Backbone Model: Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
-Backbone Robustness Clean (%) 89.1 93.9 88.1
-PGD (%) 43.9 77.1 61.8
-AutoAttack (%) 42.4 74.0 58.6
-Time Cost (s) for Defense 28 69 26
-White-Box Defense Clean (%) 95.5 95.8 93.8
-PGD (%) 84.9 91.6 85.9
-AutoAttack (%) 81.8 90.2 83.3
-Transferable Defense Victim: RaWideResNet-70-16 [35] Clean (%) 95.4 95.5 95.3
-PGD (%) 85.7 89.7 87.6
-AutoAttack (%) 82.7 88.0 85.9
-Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
-PGD (%) 72.7 78.6 79.6
-AutoAttack (%) 70.8 75.7 77.3
-  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [21] as the classifier model, and 100 steps backward propagation as Bi-Level [32].
-The transferable defenses were all evaluated using RaWideResNet-70-16 [35] and XCiT-L12 [13] as the victim models, different from the backbone models.
-Bold indicates the best performance.
-B.2 Backbone Model
-The backbone model determines the defense time cost and performance.
-As illustrated in Fig.
-8, different backbone models learn different features.
-To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec.
-1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [39].
-Figure 8: Different backbone models learn different features.
-PreActResNet-18 [19], WideResNet-94-16 [6], and Bi-Level [32] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
-As shown in Tab.
-5, the best adversarially trained model was WideResNet-94-16 [6], so it is not surprising that its white-box defense outperformed that of the other backbone models.
-However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [13]), and the time cost was much higher due to its complexity.
-We thus empirically selected PreActResNet-18 [19] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [6].
-Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
-B.3 Forward-Backward Cascade Learning
-As observed in Fig.
-9, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0).
-In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90).
-Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
-Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig.
-10.
-The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs.
-F0-B3).
-Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
-Figure 9: Forward propagation (F) converges faster while backward propagation (B) alleviates overfitting.
-The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
-The numbers in brackets represent robust accuracy against AutoAttack (Standard).
-“Alternate” represents the running of forward and backward propagation alternatively for 100 steps in total.
-Figure 10: By combining forward (F) and backward (B) propagation, Fast Preemption obtains performance comparable to that of the SOTA preemptive defense [32] by using only three steps (F1–B2).
-The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
-The numbers in brackets represent robust accuracy against AutoAttack (Standard).
-Table 6: Performance for different $N$ and optimization algorithms.
-       Ablation        Time Cost (s)        Transferable Defense
-       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
-       PGD        1        0.7        95.3        87.5        85.8
-       FGSM        1        0.04        95.4        86.8        85.4
-       PGD        20        0.8        95.2        87.7        85.9
-       FGSM        20        0.04        95.5        87.8        85.9
-  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [35] as the victim model. Bold indicates the best performance.
-B.4 Optimization Algorithm
-The results in Tab.
-6 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost.
-Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
-Table 7: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
-All proactive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
-Model Metric No Proactive Defense Bi-Level [32] Fast Preemption (ours)
-Defense Learning Backbone N/A Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
-Time Cost (s) N/A 28 0.35 0.04
-RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
-Robust 71.4 81.3 86.8 85.9
-WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
-Robust 71.2 80.9 86.4 86.0
-WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
-Robust 62.1 76.2 78.7 80.8
-WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
-Robust 64.8 77.7 81.3 83.2
-PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
-Robust 42.4 54.7 55.4 61.2
-ResNest152 [41] Clean 86.9 87.6 91.2 92.4
-Robust 62.4 73.2 75.8 78.2
-ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
-Robust 52.0 68.1 66.7 72.0
-ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
-Robust 51.4 67.5 67.8 72.8
-XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
-Robust 57.9 71.3 74.6 76.0
-Average Clean 89.9 89.8 93.4 94.1
-Robust 59.5 72.3 74.8 77.3
-  • Bold indicates the best performance.
-Appendix C Evaluation across More Models
-In the main paper, we concentrated on comparisons with top-performing adversarial training methods [6, 29].
-In this section, we extend our evaluation to include defenses tested across nine additional victim models on CIFAR-10 [26] from RobustBench [12] in a transferable context.
-These models encompass diverse architectures representative of platforms such as Twitter and Facebook, differing from both the classifier [21] and the backbone model [19].
-As shown in Tab.
-7, our method demonstrates significant performance gains compared to no proactive defense and the SOTA proactive defense [32].
-Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively.
-Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
-Appendix D Detailed Algorithms
+This supplementary material provides a detailed exposition of the algorithms underlying the proposed adversarial defense and adaptive reversion attack (Appendix A), outlines the experimental configurations for both attack and defense scenarios (Appendix B), presents evaluation results against a broader range of models (Appendix C), and offers a comprehensive analysis of visual quality (Appendix D).
+Additionally, it includes an extensive ablation study (Appendix E), examining key components such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the effectiveness of averaging multiple Fast Gradient Sign Method (FGSM) examples.
+Appendix A Detailed Algorithms
 The complete defense algorithm of Fast Preemption is presented in Algorithm 1.
@@
 5:return $\tilde{x}$
-Appendix E Experiment Settings
+Appendix B Experiment Settings
 The full experimental settings used for evaluations are listed in Tab.
-8.
+4.
 We evaluated attacks and defenses using models with various architectures, as listed in Tab.
-9.
+5.
 Our implementation partially leveraged the ART [30] and RobustBench [12] leaderboards by using PyTorch.
@@
       DeiT Base [41] (classifier)
-Appendix F Related Works
-Existing adversarial defenses are divided into three categories: training-time, test-time, and preemptive.
-Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels.
-The earliest mechanisms, e.g., adversarial training [18], incorporated various types of AEs along with their ground-truth labels in the training dataset.
-Zhang et al.
-[54] proposed TRADES, which introduces a theoretically grounded approach to balance robustness and accuracy in adversarial training by incorporating a custom loss function that controls the trade-off between clean accuracy and adversarial robustness.
-To improve the efficiency of training on large-scale datasets, Shafahi et al.
-[42] devised accelerated adversarial training in which backward-pass computations are reused.
-This entails curating a large number of specific real-world samples; Sehwag et al.
-[41] circumvented this problem by using additional data from proxy distributions learned by advanced generative models for adversarial training.
-Gowal et al.
-[19] demonstrated that generative models trained solely on the original dataset can be leveraged to artificially increase the training dataset’s size and improve adversarial robustness against $L_{p}$-norm perturbations.
-Debenedetti et al.
-[13] proposed using a custom recipe of vision transformers for adversarial training in which weaker data augmentation is used with warm-up in the augmentation intensity.
-Inspired by denoising diffusion models [23], Wang et al.
-[50] conducted adversarial training using data generated by an elucidating diffusion model [24].
-Dong et al.
-[14] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration.
-Peng et al.
-[35] devised a robust model which follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions.
-Bartoldson et al.
-[6] developed the current state-of-the-art (SOTA) robust model by scaling laws for adversarial training to assess robustness in image classifiers, revealing that even with significant computational resources, robustness asymptotes below human-level performance due to flaws in attack formulations generating invalid data, suggesting future efforts should address both computational efficiency and image validity.
-Training-time defenses have two shortcomings.
-First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness.
-Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al.
-[14], which was designed to circumvent this problem.
-Test-time defenses are divided to detection and purification.
-Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs.
-Among recent studies, Cohen et al.
-[9] devised a k-nearest neighbor model for ranking the most supportive training samples for any given validation example in training a detector.
-Zuo and Zeng [56] proposed using a binary classifier, erase-and-restore, that uses the intriguing sensitivity of $L_{2}$-norm attacks by randomly erasing certain pixels in an $L_{2}$-norm AE and then restoring them with an inpainting technique.
-Liang et al.
-[27] detected AEs by comparing the classification results of a given sample and a denoised version that was generated through adaptive noise reduction combining scalar quantization and a smoothing spatial filter.
-Most recently, Hickling et al.
-[22] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents.
-However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks [7].
-Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
-In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations.
-Earlier studies used image distortion for purification.
-For instance, Wang et al.
-[48] distorted images by randomly nullifying pixels.
-Guo et al.
-[20] preprocessed inputs with a nondifferentiable transformation.
-Raff et al.
-[37] combined a large number of individually weak defenses into a single randomized transformation to build a strong defense.
-Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious.
-Ren et al.
-[38] trained a generative model using autoencoder technology [25].
-They claimed the existence of an immune space, in which perturbations have fewer adverse effects, and they inactivated adversarial perturbations by restricting them to that subspace.
-Nie et al.
-[34] devised the DiffPure benchmark, which uses denoising diffusion models [23].
-DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process.
-The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [46].
-Lastly, Alfarra et al.
-[3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier.
-However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper.
-This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios.
-While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model.
-Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion).
-Furthermore, the approach of Alfarra et al.
-[3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
-Preemptive defenses aim to generate REs that replace original images, which are then discarded.
-These REs are more challenging to manipulate and thus resist adversarial attacks.
-Salman et al.
-[40] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples.
-They were learned using projected gradient descent (PGD) [31] and improved both in-distribution performance and robustness against unforeseen perturbations.
-Moon et al.
-[32] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy.
-However, its time cost is huge, and its transferability is low.
-Frosio and Kautz [17] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained.
-It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
+Appendix C Evaluation across More Models
+In the main paper, we concentrated on comparisons with top-performing adversarial training methods [6, 26].
+In this section, we extend our evaluation to include defenses tested across nine additional victim models on CIFAR-10 [23] from RobustBench [12] in a transferable context.
+These models encompass diverse architectures representative of platforms such as Twitter and Facebook, differing from both the classifier [21] and the backbone model [19].
+As shown in Tab.
+6, our method demonstrates significant performance gains compared to no preemptive defense and the SOTA preemptive defense [29].
+Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively.
+Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
+Table 6: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+All preemptive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
+Model Metric No Preemptive Defense Bi-Level [29] Fast Preemption (ours)
+Defense Learning Backbone N/A Bi-Level [29] WideResNet-94-16 [6] PreActResNet-18 [19]
+Time Cost (s) N/A 28 0.35 0.04
+RaWideResNet-70-16 [32] Clean 93.5 92.1 95.5 95.4
+Robust 71.4 81.3 86.8 85.9
+WideResNet-70-16 [47] Clean 93.3 91.5 95.5 95.3
+Robust 71.2 80.9 86.4 86.0
+WideResNet-34-10 [33] Clean 91.5 90.2 95.1 95.1
+Robust 62.1 76.2 78.7 80.8
+WideResNet-28-10 [49] Clean 93.4 91.5 95.4 95.6
+Robust 64.8 77.7 81.3 83.2
+PreActResNet-18 [48] Clean 84.2 87.3 90.1 91.6
+Robust 42.4 54.7 55.4 61.2
+ResNest152 [38] Clean 86.9 87.6 91.2 92.4
+Robust 62.4 73.2 75.8 78.2
+ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
+Robust 52.0 68.1 66.7 72.0
+ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
+Robust 51.4 67.5 67.8 72.8
+XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
+Robust 57.9 71.3 74.6 76.0
+Average Clean 89.9 89.8 93.4 94.1
+Robust 59.5 72.3 74.8 77.3
+  • Bold indicates the best performance.
+Appendix D Visual Quality
+The noise budgets utilized (as detailed in Tab.
+4) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection.
+As illustrated in Fig.
+7, for 2,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [46] and Learned Perceptual Image Patch Similarity (LPIPS) [52] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images.
+Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
+Figure 7: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA preemptive defense [29].
+All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images.
+SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
+D.1 Structural Similarity Index (SSIM)
+The SSIM is a metric for measuring similarity between two images.
+It considers luminance, contrast, and structural information.
+The SSIM between two images $x$ (reference) and $y$ (distorted) is defined as:
+$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (15)
+where:
+  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
+  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
+  • $\sigma_{xy}$: Covariance between $x$ and $y$.
+  • $C_{1}$ and $C_{2}$: Small constants added to stabilize the metric in the presence of weak denominators, defined as:
+    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
+    where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit images), and $K_{1}$ and $K_{2}$ are small positive constants (commonly, $K_{1}=0.01$ and $K_{2}=0.03$).
+The SSIM value ranges from -1 to 1, where:
+  • 1: Perfect similarity.
+  • 0: No similarity.
+  • Negative values indicate dissimilarity.
+SSIM is typically calculated locally using a sliding window across the image, and the overall SSIM is the average of local SSIM values.
+Higher SSIM values denote subtler deviations from the reference images.
+D.2 Learned Perceptual Image Patch Similarity (LPIPS)
+The LPIPS is a perceptual similarity measure that evaluates the perceptual distance between two images.
+It uses deep neural network features to compare the visual content and correlates better with human perception than traditional metrics like SSIM.
+The LPIPS distance between two images $x$ (reference) and $y$ (distorted) is computed as:
+$LPIPS(x,y)=$ (17)
+$\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}w_{l}\|\hat{\phi}_{l}(x)_{h,w}-\hat{\phi}_{l}(y)_{h,w}\|_{2}^{2},$
+where:
+  • $l$: Layer index in the deep network.
+  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
+  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
+  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
+  • $w_{l}$: Learned scalar weights for each layer $l$ (trained to match human judgments).
+LPIPS uses a pretrained deep neural network to extract feature maps from intermediate layers.
+The metric normalizes the feature maps and computes the difference between them, weighted according to the learned weights $w_{l}$.
+The LPIPS value typically ranges from 00 (perfect perceptual similarity) to higher values indicating greater dissimilarity.
+LPIPS provides a perceptually grounded comparison, making it especially useful for evaluating subtle changes in visual quality.
+Appendix E Ablation Studies
+We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset.
+The attacks were PGD [28] and AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
+As illustrated in Fig.
+6 of Sec.
+7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [29] in terms of clean and robust accuracies and time cost.
+These studies and results were aimed at identifying the optimal strategies.
+Table 7: Performance with different classifiers.
+       Ablation        Time Cost(s)        Transferable Defense
+       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
+       Ground Truths        100        28        97.2        86.9        83.1
+       ResNeXt        95.9        28        95.4        85.7        82.7
+       Bi-Level        89.1        29        91.9        83.9        81.0
+  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [29].
+The transferable defenses were all evaluated using RaWideResNet-70-16 [32] as the victim model.
+Bold indicates the best performance.
+E.1 Classifier
+In all preemptive defenses, correct labeling guarantees that optimization moves in the correct direction.
+As demonstrated in Tab.
+7, a classifier with higher accuracy leads to better performance.
+As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning.
+Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction.
+Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
+Table 8: Performance of different backbone models.
+Backbone Model: Bi-Level [29] WideResNet-94-16 [6] PreActResNet-18 [19]
+Backbone Robustness Clean (%) 89.1 93.9 88.1
+PGD (%) 43.9 77.1 61.8
+AutoAttack (%) 42.4 74.0 58.6
+Time Cost (s) for Defense 28 69 26
+White-Box Defense Clean (%) 95.5 95.8 93.8
+PGD (%) 84.9 91.6 85.9
+AutoAttack (%) 81.8 90.2 83.3
+Transferable Defense Victim: RaWideResNet-70-16 [32] Clean (%) 95.4 95.5 95.3
+PGD (%) 85.7 89.7 87.6
+AutoAttack (%) 82.7 88.0 85.9
+Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
+PGD (%) 72.7 78.6 79.6
+AutoAttack (%) 70.8 75.7 77.3
+  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [21] as the classifier model, and 100 steps backward propagation as Bi-Level [29].
+The transferable defenses were all evaluated using RaWideResNet-70-16 [32] and XCiT-L12 [13] as the victim models, different from the backbone models.
+Bold indicates the best performance.
+E.2 Backbone Model
+The backbone model determines the defense time cost and performance.
+As illustrated in Fig.
+8, different backbone models learn different features.
+To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec.
+1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [36].
+Figure 8: Different backbone models learn different features.
+PreActResNet-18 [19], WideResNet-94-16 [6], and Bi-Level [29] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
+As shown in Tab.
+8, the best adversarially trained model was WideResNet-94-16 [6], so it is not surprising that its white-box defense outperformed that of the other backbone models.
+However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [13]), and the time cost was much higher due to its complexity.
+We thus empirically selected PreActResNet-18 [19] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [6].
+Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
+E.3 Forward-Backward Cascade Learning
+As observed in Fig.
+9, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0).
+In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90).
+Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
+Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig.
+10.
+The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs.
+F0-B3).
+Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
+Figure 9: Forward propagation (F) converges faster while backward propagation (B) alleviates overfitting.
+The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
+The numbers in brackets represent robust accuracy against AutoAttack (Standard).
+“Alternate” represents the running of forward and backward propagation alternatively for 100 steps in total.
+Figure 10: By combining forward (F) and backward (B) propagation, Fast Preemption obtains performance comparable to that of the SOTA preemptive defense [29] by using only three steps (F1–B2).
+The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively.
+The numbers in brackets represent robust accuracy against AutoAttack (Standard).
+Table 9: Performance for different $N$ and optimization algorithms.
+       Ablation        Time Cost (s)        Transferable Defense
+       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
+       PGD        1        0.7        95.3        87.5        85.8
+       FGSM        1        0.04        95.4        86.8        85.4
+       PGD        20        0.8        95.2        87.7        85.9
+       FGSM        20        0.04        95.5        87.8        85.9
+  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [32] as the victim model. Bold indicates the best performance.
+E.4 Optimization Algorithm
+The results in Tab.
+9 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost.
+Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.